### 06 Mar 2026 v2

# Fusions of One-Variable First-Order Modal Logics

Abstract
We investigate preservation results for the independent fusion of one-variable first-order modal logics. We show that, without equality, Kripke completeness and decidability of the global and local consequence relation are preserved, under both expanding and constant domain semantics. By contrast, Kripke completeness and decidability are not preserved for fusions with equality and non-rigid constants (or, equivalently, counting up to one), again for the global and local consequence and under both expanding and constant domain semantics. This result is shown by encoding Diophantine equations. Even without equality, the finite model property is only preserved in the local case. Finally, we view fusions of one-variable modal logics as fusions of propositional modal logics sharing an S5 modality and provide a general sufficient condition for transfer of Kripke completeness and decidability (but not of finite model property).
1 Introduction
The study of combined modal logics has been an active area of research for many years [12, 25, 35]. Only a few methods of combining modal systems have proved to be robust in the sense that key properties of the component logics are always inherited by the resulting combined logic. A central example of such a robust combination is fusion: a basic method in which the component logics are combined without introducing axioms that mix their modal operators, and, correspondingly, without imposing any semantic constraints that relate the distinct accessibility relations. In fact, properties such as Kripke completeness, decidability, finite model property, and Craig interpolation are all preserved under forming the fusion of two (or more) classical propositional modal logics, for both the local and global consequence relations [31, 21, 10, 34, 4, 15, 2, 6].
The aim of this contribution is to initiate a systematic investigation into the extent to which these preservation results for fusions of propositional modal logics can be lifted to (fragments of) first-order modal logics and corresponding many-dimensional modal logics. We focus in this paper on one-variable first-order modal logics, either without equality or with equality and non-rigid constants, and on corresponding commutators and products of propositional modal logics. These logics have been studied extensively [11, 18, 24, 5, 28] as they are often computationally and semantically better behaved than full first-order modal logics. For fusions without equality our main result is as follows:
Theorem A (Equality-Free Fusion).
Kripke completeness and decidability are preserved under fusions of Kripke complete one-variable first-order modal logics, for both the local and global consequence and expanding and constant domain semantics. The finite model property is preserved for the local consequence but not for global consequence. Moreover, the finite model property fails for global consequence for all non-trivial fusions under both expanding and constant domain semantics.
The proof builds on the cactus model construction [17], extending it to one‑variable modal logic via the quasimodel technique [23]. For one‑variable first‑order modal logics with equality, the presence of equality can only have a substantial impact if one also allows non‑rigid constants or, equivalently, counting quantifiers such as $\exists^{=1}x\,P(x)\equiv\exists x\,\bigl[P(x)\wedge\forall y\,\bigl(P(y)\rightarrow(x=y)\bigr)\bigr]$, which expresses that predicate $P$ is interpreted as a singleton. In this setting, drawing on known decidability results for the one‑variable first-order extension of the modal logic of inequality (also known as the ‘elsewhere’ logic) and for one‑variable temporal logics [18, 19, 13] and by encoding Diophantine equations and Minsky machines to show undecidability, we establish the following non‑transfer result:
Theorem B (Fusion with Equality).
Decidability and recursive axiomatisability are not preserved for fusions of one-variable first-order modal logics with equality, neither for the local nor global consequence and neither for expanding nor constant domain semantics. Moreover, the global consequence is undecidable for all non-trivial fusions for constant domain semantics.
We stress that it would be a mistake to infer that fusions of one‑variable modal logics with equality are broadly undecidable. In fact, when based on standard modal systems such as polymodal K or S5, one‑variable modal logics with equality remain decidable, even in coNExpTime [1].
Since the one‑variable fragment of first‑order logic without equality can be identified with propositional modal logic S5, one‑variable first‑order modal logics can often be viewed as propositional multimodal logics equipped with an S5 modality. Working within this perspective, we provide a sufficient condition for the transfer of Kripke completeness and decidability in fusions of propositional modal logics that share an S5 modality. We say that a propositional modal logic admits $E$-homogeneous models for an equivalence relation $E$ interpreting an S5 modality $\Box_{E}$ if, for sufficiently large infinite cardinals $\kappa$, it is complete for Kripke models in which, for every EE‑equivalence class $W$ of worlds, each formula is either not satisfied anywhere in $W$ or is satisfied in $\kappa$-many worlds of $W$.
Theorem C (Propositional Fusion with Shared S5).
Kripke completeness and decidability are preserved under fusions of propositional modal logics admitting $E$-homogeneous models for a shared S5 modality $\Box_{E}$, for both the local and global consequence.
The conditions of Theorem C hold, for example, for Kripke complete (semi)commutators with S5 and for (expanding) product modal logics with S5. Hence, it provides transfer results for corresponding one-variable modal logics. We note that, in contrast to the proof of Theorem A that uses quasimodels, one cannot directly prove transfer of the finite model property based on homogeneous models. Theorem C addresses a problem posed by Baader, Ghilardi and Tinelli [2].
2 Preliminaries
We consider first-order (multi-)modal logics in languages with a single, fixed, individual variable. A language $\mathcal{L}_{M}$ contains countably many unary predicate letters, the individual variable $x$, the Booleans $\neg$ and $\land$, the quantifier symbol $\exists$, and a finite set $M$ of unary (necessity-like) modal operators, denoted by $\Box$, possibly with subscripts. Thus, in $\mathcal{L}_{M}$, atomic formulas have the form $P(x)$, where $P$ is a predicate letter. Formulas of $\mathcal{L}_{M}$, denoted by $\varphi,\psi,\ldots$, are built from atomic ones in the usual way, the only allowed quantifier being $\exists x$. The symbols $\bot$, $\to$, $\vee$, $\forall$, and $\Diamond$ are standard abbreviations. We also denote $\Box^{\leq n}\varphi=\varphi\land\Box\varphi\land\dots\land\Box^{n}\varphi$. Proposition letters $p,q,\dots$ are abbreviations for $\exists x\,P(x)$, $\exists x\,Q(x)$, etc. We also consider languages with equality: $\mathcal{L}^{=}_{M}$ contains, in addition to the symbols of $\mathcal{L}_{M}$, the binary predicate symbol == and constant symbols, denoted by $c_{1},c_{2},\dots$ Thus, in $\mathcal{L}^{=}_{M}$, a term $t$ is either $x$ or a constant $c_{i}$, and atomic formulas have the form $P(t)$ or $t_{1}=t_{2}$. We usually identify languages with sets of their formulas. The set of subformulas of a formula $\varphi$ is denoted by $\mathop{\mathit{sub}}\varphi$.
By a one-variable (first-order) $M$-logic, we mean a set of $\mathcal{L}_{M}$-formulas that includes the one-variable fragment of the classical first-order logic $\mathbf{FO}$ and formulas $\Box\bigl(P(x)\to Q(x)\bigr)\to\bigl(\Box P(x)\to\Box Q(x)\bigr)$, for all $\Box\in M$, and is closed under Modus Ponens (MP), Generalisation ($\varphi/\forall x\,\varphi$), Necessitation ($\varphi/\Box\varphi$, for each $\Box\in M$), and Substitution (of arbirary one-variable formulas for formulas of the form P(x)P(x)). [^1] One-variable fragments of first-order modal logics are one-variable logics in our sense, but not every one-variable logic is a one-variable fragment of a first-order modal logic; see, e.g., [29, Section 3]. If $L$ is a one-variable $M$-logic and $\Gamma\subseteq\mathcal{L}_{M}$, then $L+\Gamma$ denotes the smallest one-variable $M$-logic including $L\cup\Gamma$; we write $L+\varphi$ for $L+\{\varphi\}$. We note that every one-variable $M$-logic contains the converse Barcan formula $\textit{CBF}_{\Box}=\Box\forall x\,P(x)\to\forall x\,\Box P(x)$, for each $\Box\in M$.
Let $L$ be a one-variable $M$-logic and $\varphi,\psi\in\mathcal{L}_{M}$. We write $\varphi\vdash_{L}\psi$ and say that $\psi$ is a local consequence of $\varphi$ in $L$ if $\psi$ belongs to the smallest set including $L\cup\{\varphi\}$ and closed under MP and Generalization. By the deduction theorem, $L$ is decidable iff $\vdash_{L}$ is decidable. We write $\varphi\vdash_{L}^{\ast}\psi$ and say that $\psi$ is a global consequence of $\varphi$ in $L$ if $\psi$ belongs to the smallest set including $L\cup\{\varphi\}$ and closed under MP, Generalization, and Necessitation for all $\Box\in M$. We call $L$ globally decidable if the global consequence in $L$ is decidable.
The fusion of a one-variable $M_{1}$-logic $L_{1}$ and a one-variable $M_{2}$-logic $L_{2}$, where $M_{1}\cap M_{2}=\emptyset$, is the smallest one-variable $(M_{1}\cup M_{2})$-logic $L_{1}\otimes L_{2}$ that includes $L_{1}\cup L_{2}$.
We are mostly concerned with logics defined in terms of Kripke semantics. A Kripke $M$-frame is a tuple $\mathfrak{F}=(W,\{R_{\Box}\}_{\Box\in M})$, where $W$ is a non-empty set of worlds and $R_{\Box}\subseteq W\times W$, for all $\Box\in M$. We write $R_{\Box}(w)$ for $\{w^{\prime}\in W\mid wR_{\Box}w^{\prime}\}$. We say that $\mathfrak{F}$ is non-trivial if there are distinct $w,v\in W$ with $wR_{\Box}v$, for some $\Box\in M$. A Kripke cd-model based on $\mathfrak{F}$ is a pair $\mathfrak{M}=(\mathfrak{F},I)$, where $I=\{\langle\Delta,I_{w}\rangle\mid w\in W\}$ is a family of classical first-order models. A Kripke $\mathit{xd}$-model based on $\mathfrak{F}$ is a pair $\mathfrak{M}=(\mathfrak{F},I)$, where $I=\{\langle\Delta_{w},I_{w}\rangle\mid w\in W\}$ is a family of classical first-order models satisfying the expanding domains condition: $\Delta_{w}\subseteq\Delta_{v}$, for all $w,v\in W$ with $wR_{\Box}v$ and all $\Box\in M$. For languages with equality and constants, the classical models associated with worlds are required to be normal (i.e., == is interpreted as identity); the resultant Kripke models are called Kripke $d$-models with equality, for $d\in\{\textit{xd},\textit{cd}\}$; note that rigid interpretation of constants, i.e., $I_{w}(c_{k})=I_{v}(c_{k})$, for all $w,v\in W$, is not assumed. We say that a Kripke model is finite if its set of worlds is finite and every classical model associated with its worlds is finite.
Since cd-models are a special case of xd-models, we define truth and validity only for the latter. Let $\mathfrak{M}=(W,\{R_{\Box}\}_{\Box\in M},\langle\Delta_{w},I_{w}\rangle_{w\in W})$ be an xd-model. An assignment in $\mathfrak{M}$ is a map $\alpha$ sending $x$ to an element of the set $\bigcup_{w\in W}\Delta_{w}$. The truth of a one-variable $\mathcal{L}_{M}$-formula at world $w$ of $\mathfrak{M}$ under assignment $\alpha$ is defined by recursion (the value of the term $t$ at $w$ under $\alpha$ is denoted by $|t|_{w}^{\alpha}$):
  • $\mathfrak{M},w\models^{\alpha}P(t)$ if $|t|_{w}^{\alpha}\in I_{w}(P)$;
  • $\mathfrak{M},w\models^{\alpha}\neg\varphi$ if $\mathfrak{M},w\not\models^{\alpha}\varphi$;
  • $\mathfrak{M},w\models^{\alpha}\varphi_{1}\land\varphi_{2}$ if $\mathfrak{M},w\models^{\alpha}\varphi_{1}$ and $\mathfrak{M},w\models^{\alpha}\varphi_{2}$;
  • $\mathfrak{M},w\models^{\alpha}\Box\varphi$ if $\mathfrak{M},w^{\prime}\models^{\alpha}\varphi$, for every $w^{\prime}\in R_{\Box}(w)$;
  • $\mathfrak{M},w\models^{\alpha}\exists x\,\varphi$ if $\mathfrak{M},w\models^{\alpha^{\prime}}\varphi$, for some $\alpha^{\prime}$ such that $\alpha^{\prime}(x)\in\Delta_{w}$.
In languages $\mathcal{L}_{M}^{=}$ with equality, additionally,
  • $\mathfrak{M},w\models^{\alpha}t_{1}=t_{2}$ if $|t_{1}|_{w}^{\alpha}=|t_{2}|_{w}^{\alpha}$.
We say that $\varphi$ is true at $w\in W$, and write $\mathfrak{M},w\models\varphi$, if $\mathfrak{M},w\models^{\alpha}\varphi$ holds for every $\alpha$ with $\alpha(x)\in\Delta_{w}$, and say that $\varphi$ true in $\mathfrak{M}$, and write $\mathfrak{M}\models\varphi$, if $\mathfrak{M},w\models\varphi$ holds for every $w\in W$.
Let $d\in\{\mathit{xd},\mathit{cd}\}$. We write $\mathfrak{F}\models_{d}\varphi$ if $\varphi$ is true in every $d$-model based on $\mathfrak{F}$. Then $\mathfrak{F}$ is a $d$-frame for a logic $L$, in symbols $\mathfrak{F}\models_{d}L$, if $\mathfrak{F}\models_{d}\varphi$ for all $\varphi\in L$. An $\mathcal{L}_{M}$-formula $\varphi$ is $d$-valid on a class $\mathscr{C}$ of Kripke $M$-frames (written as $\mathscr{C}\models_{d}\varphi$) if $\mathfrak{F}\models_{d}\varphi$, for all $\mathfrak{F}\in\mathscr{C}$. Analogous notions are defined for formulas and models with equality. We define the following sets of formulas:
[…]
These sets are one-variable $M$-logics. Note that, for every $\mathscr{C}$, we have $\mathscr{C}\models_{\textit{cd}}\forall x\,\Box P(x)\to\Box\forall x\,P(x)$ (the Barcan formula, $\textit{BF}_{\Box}$), for all $\Box\in M$.
Let $L$ be a one-variable $M$-logic, $\mathcal{C}$ a class of $M$-frames, and $d\in\{\textit{xd},\textit{cd}\}$. We say $L$ is $d$-complete for $\mathcal{C}$ if $L=\mathop{\mathit{Log}}_{d}\mathscr{C}$, and that $L$ is globally $d$-complete for $\mathcal{C}$ in case $\varphi\vdash_{L}^{*}\psi$ iff $\mathfrak{M}\models\varphi$ implies $\mathfrak{M}\models\psi$, for every $d$-model $\mathfrak{M}$ based on a frame from $\mathcal{C}$. Note that global $d$-completeness implies $d$-completeness, but the converse is not always true. We say that $L$ has the finite $d$-model property ($d$-fmp) if, for every $\varphi\notin L$, there exists a finite $d$-model $\mathfrak{M}$ such that $\mathfrak{M}\models L$ and $\mathfrak{M}\not\models\varphi$. We say that $L$ has the global finite $d$-model property (global $d$-fmp) if, for every $\varphi$ and $\psi$ with $\varphi\not\vdash^{*}_{L}\psi$, there exists a finite $d$-model $\mathfrak{M}$ such that $\mathfrak{M}\models L\cup\{\varphi\}$ and $\mathfrak{M}\not\models\psi$.
The fusion of Kripke frames $\mathfrak{F}_{1}=(W,\{R_{\Box}\}_{\Box\in M_{1}})$ and $\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{2}})$, where $M_{1}\cap M_{2}=\emptyset$, is the frame $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}=(W,\{R_{\Box}\}_{\Box\in M_{1}\cup M_{2}})$. If $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ are classes of Kripke frames, then $\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ denotes the class of frames comprising $\mathfrak{F}_{1}\otimes\mathfrak{F}_{2}$, for all $\mathfrak{F}_{1}\in\mathscr{C}_{1}$ and $\mathfrak{F}_{2}\in\mathscr{C}_{2}$ with the same set of worlds.
We shall also be concerned with propositional modal logics closely related to one-variable first-order logics. Propositional formulas are built out of proposition letters $p,q,\ldots$ using the Booleans and modalities from $M$. A propositional $M$-logic is a set of propositional $M$-formulas containing the classical propositional logic, formulas $\Box(p\to q)\to(\Box p\to\Box q)$, for all $\Box\in M$, and closed under MP, Necessitation for all $\Box\in M$, and Propositional Substitution. If $L$ is a propositional $M$-logic and $\Gamma$ a set of propositional $M$-formulas, then $L+\Gamma$ denotes the smallest propositional $M$-logic including $L\cup\Gamma$; we write $L+\varphi$ for $L\cup\{\varphi\}$. A (propositional) Kripke model based on $\mathfrak{F}$ is a pair $\mathfrak{M}=(\mathfrak{F},V)$ with $V(p)\subseteq W$ for all propositional letters $p$. The truth of a formula at world $w$ of $\mathfrak{M}$ under assignment $V$ is defined by recursion as usual:
  • $\mathfrak{M},w\models p$ if $w\in V(p)$;
  • $\mathfrak{M},w\models\neg\varphi$ if $\mathfrak{M},w\not\models\varphi$;
  • $\mathfrak{M},w\models\varphi_{1}\land\varphi_{2}$ if $\mathfrak{M},w\models\varphi_{1}$ and $\mathfrak{M},w\models\varphi_{2}$;
  • $\mathfrak{M},w\models\Box\varphi$ if $\mathfrak{M},w^{\prime}\models\varphi$, for every $w^{\prime}\in R_{\Box}(w)$;
We write $\mathfrak{F}\models\varphi$ if $\varphi$ is true in every world in $\mathfrak{F}$ for all models $\mathfrak{M}$ based on $\mathfrak{F}$. Then $\mathfrak{F}$ is a frame for a logic $L$, in symbols $\mathfrak{F}\models L$, if $\mathfrak{F}\models\varphi$ for all $\varphi\in L$. The logic $L$ is Kripke complete if $\varphi\in L$ whenever $\mathfrak{F}\models\varphi$ for all frames $\mathfrak{F}$ for $L$. Global Kripke completeness and (global) finite model property are defined analogously to the first-order case.
The fusion of a propositional $M_{1}$-logic $L_{1}$ and a propositional $M_{2}$-logic $L_{2}$, where $M_{1}\cap M_{2}=\emptyset$, is smallest propositional $(M_{1}\cup M_{2})$-logic $L_{1}\otimes L_{2}$ that includes $L_{1}\cup L_{2}$.
The semicommutator [^2] In general, semicommutators are also required to contain confluence axioms for commuting modalities; these are derivable in semicommutators with $\mathbf{S5}$. and the commutator  [23, 25] of a propositional $M$-logic $L$ with the propositional monomodal logic $\mathbf{S5}$ are the propositional $(M\cup\{\Box_{E}\})$-logics
$L,\textbf{S5}]^{\textit{xd}}\ \ =\ \ L\otimes\textbf{S5}\ +\ \bigl\{(\textit{lcom}_{\Box})\mid\Box\in M\bigr\}\quad\text{ and }\quad[L,\textbf{S5}]^{\textit{cd}}\ \ =\ \ [L,\textbf{S5}]^{\textit{xd}}\ +\ \bigl\{(\textit{rcom}_{\Box})\mid\Box\in M\bigr\},$
where $(\textit{lcom}_{\Box})$ and $(\textit{rcom}_{\Box})$ are the following formulas with their Kripke frame correspondents:
$\Box\Box_{E}\,p\to\Box_{E}\Box\,p,$ $E\circ R\subseteq R\circ E,$ ($\textit{lcom}_{\Box}$)
$\Box_{E}\Box\,p\to\Box\Box_{E}\,p,$ $R\circ E\subseteq E\circ R.$ ($\textit{rcom}_{\Box}$)
There exists a well-know translation of one-variable $M$-formulas into propositional $(M\cup\{\Box_{E}\})$-formulas: $P(x)^{\ast}=p$; $(\neg\varphi)^{\ast}=\neg\varphi^{\ast}$; $(\varphi_{1}\land\varphi_{2})^{\ast}=\varphi_{1}^{\ast}\land\varphi_{2}^{\ast}$; $(\Box\varphi)^{\ast}=\Box\varphi^{\ast}$, for all $\Box\in M$; $(\exists x\,\varphi)^{\ast}=\Diamond_{E}\varphi^{\ast}$. We note that $\textit{CBF}_{\Box}^{\ast}=(\textit{lcom}_{\Box})$ and $\textit{BF}_{\Box}^{\ast}=(\textit{rcom}_{\Box})$. If $\Gamma\subseteq\mathcal{L}_{M}$, then $\Gamma^{\ast}=\{\varphi^{\ast}\mid\varphi\in\Gamma\}$.
If $L$ is a propositional $M$-logic, then the smallest one-variable $M$-logic containing all the substitution instances of formulas from $L$ (here, we substitute arbitrary one-variable formulas for proposition letters) is denoted by $\mathcal{Q}L$. The following is proven in Appendix A:
Lemma 2.1.
If $L$ is a propositional $M$-logic, then $\mathcal{Q}L^{\ast}=[L,\mathbf{S5}]^{\mathit{xd}}$ and $(\mathcal{Q}L+\mathit{BF})^{\ast}=[L,\mathbf{S5}]^{\mathit{cd}}$.
3 Fusions of Equality-Free One-Variable Logics
Theorem 3.1.
Let $\mathcal{C}_{1},\mathcal{C}_{2}$ be non-empty classes of frames closed under disjoint unions, $d\in\{\textit{xd},\textit{cd}\}$, and $L_{i}$ be one-variable modal logics (globally) $d$-complete for $\mathcal{C}_{i}$, for both $i=1,2$. Then
  • $L_{1}\otimes L_{2}$ is (respectively, globally) $d$-complete for $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$;
  • $L_{1}\otimes L_{2}$ is (globally) decidable if $L_{1}$ and $L_{2}$ are both (globally) decidable.
The proof of the transfer results is based on the cactus model construction [17] lifted to the one-variable modal logic using the quasimodel technique; see, e.g., [23]. As has already been observed in the propositional setting [21, 2], the analysis of local consequence is substantially more intricate than that of global consequence. In the sequel, we denote the language of $L_{1}\otimes L_{2}$ by $\mathcal{L}$ and the language of $L_{i}$ by $\mathcal{L}_{i}$, for $i=1,2$. For simplicity, we assume that $L_{1}$ and $L_{2}$ are monomodal logics with modalities, respectively, $\Box_{1}$ and $\Box_{2}$. We require two main ingredients: (a) types and quasistates for a finite representation of any number of domain elements and (b) surrogates for separating the two components of the fusion.
Let $\Phi$ be a finite set of $\mathcal{L}$-formulas closed under subformulas. A $\Phi$-type $t$ is a maximal Boolean-consistent subset of $\{\psi,\neg\psi\mid\psi\in\Phi\}$, that is,
$\neg\psi\in t\quad$ $\text{ iff }\quad\psi\notin t,$ $\text{for each }\psi\in\Phi,$ (t1)
$\psi_{1}\land\psi_{2}\in t\quad$ $\text{ iff }\quad\psi_{1},\psi_{2}\in t,$ $\text{for each }\psi_{1}\land\psi_{2}\in\Phi.$ (t2)
The number of $\Phi$-types is bounded by $2^{|\Phi|}$ (for $\neg\psi\in\Phi$, a type contains either $\neg\psi$ or both $\neg\neg\psi$ and $\psi$). A $\Phi$-quasistate is a non-empty set $\boldsymbol{q}$ of $\Phi$-types such that
[…]
With a $\Phi$-quasistate $\boldsymbol{q}$, we associate the following realisability $\mathcal{L}$-sentence:
$\hat{\boldsymbol{q}}\ \ =\ \ \forall x\,\bigvee_{t\in\boldsymbol{q}}t(x)\ \ \land\ \ \bigwedge_{t\in\boldsymbol{q}}\exists x\,t(x),$
where $t(x)$ denotes the conjunction of formulas in $t$. If $\boldsymbol{Q}$ is a set of $\Phi$-quasistates, then $\hat{\boldsymbol{Q}}=\bigvee\{\hat{\boldsymbol{q}}\mid\boldsymbol{q}\in\boldsymbol{Q}\}$.
Let $\varphi$ be an $\mathcal{L}$-formula. For each $\Box_{i}\psi\in\mathop{\textit{sub}}\varphi$, we introduce a fresh unary predicate $P_{\Box_{i}\psi}$ and define the surrogate of $\Box_{i}\psi$ by taking $P_{\Box_{i}\psi}(x)$ if $\psi$ has a free variable and $\forall x\,P_{\Box_{i}\psi}(x)$ if $\psi$ is closed. Denote by $\varphi^{\mathbf{1}}$ the $\mathcal{L}_{1}$-formula obtained from $\varphi$ by replacing $\Box_{2}\psi$ not in the scope of another $\Box_{2}$ with its surrogate; symmetrically, $\varphi^{\mathbf{2}}$ denotes the $\mathcal{L}_{2}$-formula obtained from $\varphi$ by replacing its maximal $\Box_{1}\psi$-subformulas with their surrogates.
3.1 Global Consequence
To establish xd-completeness of the global consequence claimed in Theorem 3.1, we show the following lemma. Global consequence is then decidable by a simple enumeration of all possible $\Phi$-quasistates.
Lemma 3.2.
Let $\varphi$ and $\psi$ be $\mathcal{L}$-formulas. Denote $\Phi=\mathop{\textit{sub}}\varphi\cup\mathop{\textit{sub}}\psi$. Then the following are equivalent:
  • (i)
    $\varphi\not\vdash_{L_{1}\otimes L_{2}}^{*}\psi$;
  • (ii)
    there is an xd-model $\mathfrak{M}$ based on a frame from $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$ such that $\mathfrak{M}\models\varphi$ but $\mathfrak{M}\not\models\psi$;
  • (iii)
    there is a set $\boldsymbol{Q}$ of $\Phi$-quasistates such that
    • (C1)
      $\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}\not\vdash_{L_{1}}^{*}\psi^{\mathbf{1}}$,
    • (C2)
      $\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}\not\vdash_{L_{1}}^{*}\neg\hat{\boldsymbol{q}}_{i}^{\mathbf{1}}$, for each $\boldsymbol{q}_{i}\in\boldsymbol{Q}$,
    • (C3)
      $\smash{\hat{\boldsymbol{Q}}}^{\mathbf{2}}\not\vdash_{L_{2}}^{*}\neg\hat{\boldsymbol{q}}_{i}^{\mathbf{2}}$, for each $\boldsymbol{q}_{i}\in\boldsymbol{Q}$.
Proof.
Implication $(\textit{ii})\Rightarrow(\textit{i})$ follows form soundness of $\vdash^{*}_{L_{1}\otimes L_{2}}$.
$(\textit{i})\Rightarrow(\textit{iii})$ Suppose $\varphi\not\vdash_{L_{1}\otimes L_{2}}^{*}\psi$ and consider the following set $\boldsymbol{Q}$ of $\Phi$-quasistates:
$\boldsymbol{Q}=\bigl\{\boldsymbol{q}\text{ a $\Phi$-quasistate }\mid\varphi\not\vdash^{*}_{L_{1}\otimes L_{2}}\neg\hat{\boldsymbol{q}}\bigr\}$.
Clearly, $\bigvee_{\boldsymbol{q}}\hat{\boldsymbol{q}}$, for all $\Phi$-quasistates $\boldsymbol{q}$, is an $L_{1}\otimes L_{2}$-tautology, and so $\varphi\vdash^{*}_{L_{1}\otimes L_{2}}\bigvee_{\boldsymbol{q}}\hat{\boldsymbol{q}}$. Observe that $\hat{\boldsymbol{Q}}$ consists of the disjuncts left after removing all $\boldsymbol{q}$ with $\varphi\vdash^{*}_{L_{1}\otimes L_{2}}\neg\hat{\boldsymbol{q}}$. Thus, we obtain $\varphi\vdash^{*}_{L_{1}\otimes L_{2}}\hat{\boldsymbol{Q}}$. Next, $\varphi\not\vdash^{*}_{L_{1}\otimes L_{2}}\psi$ implies $\varphi\land\hat{\boldsymbol{Q}}\not\vdash^{*}_{L_{1}\otimes L_{2}}\psi$. Moreover, by definition, we also have $\hat{\boldsymbol{Q}}\not\vdash^{*}_{L_{1}\otimes L_{2}}\neg\hat{\boldsymbol{q}}_{i}$ and therefore $\varphi\land\hat{\boldsymbol{Q}}\not\vdash^{*}_{L_{1}\otimes L_{2}}\neg\hat{\boldsymbol{q}}_{i}$, for every $\boldsymbol{q}_{i}\in\boldsymbol{Q}$. Then (C1)(C3) follow by contraposition: for example, if we assume $\varphi\land\hat{\boldsymbol{Q}}\not\vdash^{*}_{L_{1}\otimes L_{2}}\psi$ but $\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}\vdash_{L_{1}}^{*}\psi^{\mathbf{1}}$, then the latter also provides an inference in $L_{1}\otimes L_{2}$, whence, as a substitution instance, we obtain $\varphi\land\hat{\boldsymbol{Q}}\vdash^{*}_{L_{1}\otimes L_{2}}\psi$ contrary to our assumption.
$(\textit{iii})\Rightarrow(\textit{ii})$ Suppose there is a set $\boldsymbol{Q}$ of $\Phi$-quasistates such that (C1)(C3) hold. To describe the construction of the witnessing xd-model $\mathfrak{M}$, we require the following set of definitions.
Definition 3.3.
A $\Phi$-run $r$ through $(W,R_{i})$ is a function from an $R_{i}$-closed subset of $W$, denoted by $\mathop{\mathsf{dom}}r$, to $\Phi$-types. For $w\in\mathop{\mathsf{dom}}r$, a run $r$ is called
  • $i$-coherent at $w$ if $\Box_{i}\psi\in r(w)$ implies $\psi\in r(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$;
  • $i$-saturated at $w$ if $\Box_{i}\psi\in r(w)$ whenever $\Box_{i}\psi\in\Phi$ and $\psi\in r(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$.
We say that $r$ is $i$-coherent on $W^{\prime}\subseteq W$ if $r$ is $i$-coherent at each $w\in\mathop{\mathsf{dom}}r\cap W^{\prime}$ and $i$-coherent if $r$ is $i$-coherent on $W$; similarly for $r$ being $i$-saturated.
An $i$-quasimodel $\mathfrak{Q}$ for $\Phi$ is a tuple of the form $((W,R_{i}),w_{0},\mathfrak{R})$, where $(W,R_{i})\in\mathcal{C}_{i}$, $w_{0}\in W$ and $\mathfrak{R}$ is a set of $\Phi$-runs through $(W,R_{i})$ such that each $r\in\mathfrak{R}$ is $i$-coherent and $i$-saturated, and each $\mathfrak{R}(w)$ is a $\Phi$-quasistate, for $w\in W$, where we denote $\mathfrak{R}_{w}=\{r\in\mathfrak{R}\mid w\in\mathop{\mathsf{dom}}r\}$ and $\mathfrak{R}(w)=\{r(w)\mid r\in\mathfrak{R}_{w}\}$.
Definition 3.4.
An $i$-cactus $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ for $\Phi$ with thorns $V$ is defined inductively.
basis
A $1$-quasimodel $((W^{0},R^{0}_{1}),w_{0},\mathfrak{R}^{0})$ for $\Phi$ gives rise to a $2$-cactus $((W^{0},R_{1}^{0},\emptyset),w_{0},\mathfrak{R}^{0})$ for $\Phi$ with thorns $W^{0}$. Note that each $r\in\mathfrak{R}^{0}$ is vacuously $2$-coherent, but not necessarily $2$-saturated. A $1$-cactus is defined symmetrically.
induction step
Let $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ be a $2$-cactus for $\Phi$ with thorns $V\subseteq W$. Let $\{\mathfrak{Q}^{v}\mid v\in V\}$ be a collection of $2$-quasimodels $\mathfrak{Q}^{v}=((W^{v},R^{v}_{2}),v,\mathfrak{R}^{v})$ for $\Phi$ such that, for each $v\in V$, we have
$W\cap W^{v}=\{v\},\qquad W^{v}\cap W^{v^{\prime}}=\emptyset,\text{ for all }v^{\prime}\in V\text{ with }v^{\prime}\neq v,\qquad\text{ and }\qquad\mathfrak{R}(v)=\mathfrak{R}^{v}(v).$
Then $((W^{\prime},R_{1},R_{2}^{\prime}),w_{0},\mathfrak{R}^{\prime})$ is a $1$-cactus for $\Phi$ with thorns $W^{\prime}\setminus W$, where
$W^{\prime}\ =\ W\ \cup\ \bigcup\nolimits_{v\in V}W^{v},\qquad R_{2}^{\prime}\ =\ R_{2}\ \cup\ \bigcup\nolimits_{v\in V}R_{2}^{v},\\ \mathfrak{R}^{\prime}\ \ =\ \ \bigcup_{r\in\mathfrak{R}}\bigl\{r\cup\hskip-11.99998pt\bigcup_{v\in\mathop{\mathsf{dom}}r\cap V}\hskip-11.99998ptr^{v}\mid r^{v}\in(\mathfrak{R}^{v})_{v}\text{ with }r(v)=r^{v}(v),\text{ for }v\in\mathop{\mathsf{dom}}r\cap V\bigr\}\ \ \cup\ \ \bigcup_{v\in V}(\mathfrak{R}^{v}\setminus(\mathfrak{R}^{v})_{v}),$
and $f\cup\bigcup_{v\in V^{\prime}}f^{v}$ denotes the partial function that coincides with $f$ on $\mathop{\mathsf{dom}}f$ and with $f^{v}$ on $\mathop{\mathsf{dom}}f^{v}$, for each $v\in V^{\prime}$; clearly, the $\Phi$-runs in $\mathfrak{R}^{\prime}$ are well-defined. It can be seen that each $r^{\prime}\in\mathfrak{R}^{\prime}$ is $1$- and $2$-coherent as well as $2$-saturated on $W^{\prime}$, and is $1$-saturated on $W$ but not necessarily on $W^{\prime}\setminus W$.
A limit $1$-cactus for $\Phi$ is obtained as the limit of the sequence of cacti starting from a $1$-quasimodel.
We construct a limit 1-cactus for $\Phi$ by using the following procedure. We first observe that, by (C2) and global Kripke completeness of $L_{1}$, we have $\varphi^{\mathbf{1}}\in t^{\mathbf{1}}$ and so $\varphi\in t$ and thus $\varphi^{\mathbf{2}}\in t^{\mathbf{2}}$, for every $t\in\boldsymbol{q}$ and $\boldsymbol{q}\in\boldsymbol{Q}$. We then begin, by (C1) and global Kripke completeness of $L_{1}$, with an xd-model $\mathfrak{M}^{0}$ based on $(W^{0},R^{0}_{1})\in\mathcal{C}_{1}$ such that $\mathfrak{M}^{0},w\models\varphi^{\mathbf{1}}$ and $\mathfrak{M}^{0},w\models\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}$, for all $w\in W^{0}$, but $\mathfrak{M}^{0},w_{0}\not\models\psi^{\mathbf{1}}$, for some $w_{0}\in W^{0}$. With each $w\in W^{0}$ and $e\in\Delta^{0}_{w}$, we associate a uniquely determined $\Phi$-type $t_{w,e}$ such that
$\mathfrak{M}^{0},w\models t_{w,e}^{\mathbf{1}}[e].$
Let $\mathfrak{R}^{0}$ be the set of functions $r_{e}$, for $e\in\bigcup_{w\in W^{0}}\Delta^{0}_{w}$, that map each $w\in W^{0}$ with $e\in\Delta^{0}_{w}$ to the type $t_{w,e}$. It follows that each $r_{e}$ is a $1$-coherent and $1$-saturated $\Phi$-run. The $1$-quasimodel $((W^{0},R^{0}_{1}),w_{0},\mathfrak{R}^{0})$ for $\Phi$ gives rise to our initial $2$-cactus for $\Phi$ with thorns $W^{0}$. Observe that, for every $\chi\in\Phi$ and $i=1,2$, we have
$\mathfrak{M}^{0},w\models\chi^{\mathbf{i}}[e]\qquad\text{ iff }\qquad\chi^{\mathbf{i}}\in t_{w,e}^{\mathbf{i}},\qquad\text{ for all }w\in W^{0}\text{ and }e\in\Delta^{0}_{w}.$
Let $\mathfrak{C}=((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ be a $2$-cactus for $\Phi$ with thorns $V\subseteq W$. Take any $v\in V$ and consider the $\Phi$-quasistate $\mathfrak{R}(v)=\boldsymbol{q}^{v}\in\boldsymbol{Q}$ selected for $v$. By (C3) and global Kripke completeness of $L_{2}$, there is an xd-model $\mathfrak{M}^{v}$ based on $(W^{v},R_{2}^{v})\in\mathcal{C}_{2}$ such that $\mathfrak{M}^{v},v\models(\hat{\boldsymbol{q}}^{v})^{\mathbf{2}}$ and $\mathfrak{M}^{v},w\models\smash{\hat{\boldsymbol{Q}}}^{\mathbf{2}}$, for all $w\in W^{v}$. Similarly to the previous step, with each $w\in W^{v}$ and $e\in\Delta_{w}^{v}$, we associate a uniquely determined $\Phi$-type $t^{v}_{w,e}$ such that
$\mathfrak{M}^{v},w\models(t^{v}_{w,e})^{\mathbf{2}}[e].$
Let $\mathfrak{R}^{v}$ be the set of functions $r^{v}_{e}$, for $e\in\bigcup_{w\in W^{v}}\Delta_{w}^{v}$, that map each $w\in W^{v}$ with $e\in\Delta^{v}_{w}$ to the type $t^{v}_{w,e}$. We have $\mathfrak{R}^{v}(v)=\boldsymbol{q}^{v}=\mathfrak{R}(v)$. As before, $\mathfrak{Q}^{v}=((W^{v},R_{2}^{v}),v,\mathfrak{R}^{v})$, for each thorn $v\in V$, is a $2$-quasimodel for $\Phi$, so we extend the 2-cactus $\mathfrak{C}$ to a $1$-cactus $\mathfrak{C}^{\prime}$ for $\Phi$ by grafting 2-quasimodels $\mathfrak{Q}^{v}$ onto $\mathfrak{C}$ (we can assume that the $W^{v}$ are pairwise disjoint and each intersects $W$ on $\{v\}$ only).
By repeating this procedure, we obtain a limit 1-cactus $\mathfrak{C}=((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ for $\Phi$. Since the $\mathfrak{M}^{v}$ realise the respective quasimodels, for every $\chi\in\Phi$, $i=1,2$ and $v\in W$, we have
$\mathfrak{M}^{v},w\models\chi^{\mathbf{i}}[e]\ \qquad\text{ iff }\qquad\chi^{\mathbf{i}}\in(r_{e}(w))^{\mathbf{i}},\qquad\text{ for all }w\in W^{v}\text{ and }e\in\Delta^{v}_{w}.$
We now construct a Kripke model $\mathfrak{M}$ from the limit 1-cactus $\mathfrak{C}$ for $\Phi$. The frame is taken from $\mathfrak{C}$ (by assumption, both $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ are closed under disjoint unions), the domain $\Delta_{w}=\mathfrak{R}_{w}$, for each $w\in W$, and $\mathfrak{M},w\models P[r]$ iff $P(x)\in r(w)$, for every unary predicate $P$, $w\in W$ and $r\in\mathfrak{R}_{w}$. By induction on the structure of $\chi\in\Phi$, we show that, for each $w\in W$ and $r\in\mathfrak{R}_{w}$,
$\mathfrak{M},w\models\chi[r]\qquad\text{ iff }\qquad\chi^{\mathbf{1}}\in(r(w))^{\mathbf{1}}\qquad\text{ iff }\qquad\chi^{\mathbf{2}}\in(r(w))^{\mathbf{2}}.$
The basis of induction is by definition. The cases of the Boolean connectives and quantifiers follow from the definition of quasistates. The cases of the modalities follow from the fact that the $\Phi$-runs in $\mathfrak{R}$ are $i$-coherent and $i$-saturated. Thus, $\mathfrak{M},w\models\varphi$, for all $w\in W$, but $\mathfrak{M},w_{0}\not\models\psi$. ∎
The claims in Theorem 3.1 for cd-models are obtained by a straightforward modification of the construction (the runs are required to be total functions on the sets of worlds).
Theorem 3.5.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be classes of frames closed under disjoint unions containing non-trivial frames, $d\in\{\textit{xd},\textit{cd}\}$, and $L_{i}=\mathop{\mathit{Log}_{d}}\mathcal{C}_{i}$, for both $i=1,2$. Then $L_{1}\otimes L_{2}$ does not enjoy the global $d$-fmp.
Proof.
Let $\mathfrak{F}_{1}\in\mathcal{C}_{1}$ and $\mathfrak{F}_{2}\in\mathcal{C}_{2}$ be non-trivial frames, and let $\mathfrak{F}=(W,R_{1},R_{2})$ be a limit cactus frame built from disjoint copies of $\mathfrak{F}_{1}$ and $\mathfrak{F}_{2}$, and $w_{0}\in W$. Clearly, $\mathfrak{F}\in\mathcal{C}_{1}\otimes\mathcal{C}_{2}$. For $k\in\mathbb{N}$, we say that $w\in W$ is $k$ steps away from $w_{0}$ if $\mathfrak{F}$ contains a path $w^{0}R_{1}v^{0}R_{2}w^{1}R_{1}v^{1}R_{2}\cdots R_{1}v^{k-1}R_{2}w^{k}$ with $w_{0}=w^{0}$ and $w^{k}=w$, and denote by $W^{k}$ the set of worlds that are $k$ steps away from $w_{0}$. Define
$\mbox{$\Box$}\hskip-7.7778pt\hskip 1.00006pt\mbox{$\Box$}\psi=\big(p\to\Box_{1}(\neg p\to\Box_{2}(p\to\psi))\big)\wedge\big(\neg p\to\Box_{1}(p\to\Box_{2}(\neg p\to\psi))\big)\quad\text{ and }\quad\mbox{$\Diamond$}\hskip-6.66669pt\hskip 1.00006pt\mbox{$\Diamond$}\psi=\neg\mbox{$\Box$}\hskip-7.7778pt\hskip 1.00006pt\mbox{$\Box$}\neg\psi,$
for a fresh propositional letter $p$, and consider the formula
$\varphi\ \ =\ \ \forall x\,\mbox{$\Diamond$}\hskip-6.66669pt\hskip 1.00006pt\mbox{$\Diamond$}Q(x)\ \ \wedge\ \ \bigl(\exists x\,Q(x)\to\exists x\,\neg Q(x)\bigr)\ \ \wedge\ \ \forall x\,\bigl(Q(x)\to\mbox{$\Box$}\hskip-7.7778pt\hskip 1.00006pt\mbox{$\Box$}Q(x)\bigr).$
Define a model $\mathfrak{M}_{0}=(\mathfrak{F},\langle\mathbb{N},I_{w}\rangle_{w\in W})$ by setting $I_{w}(Q)=\{1,\ldots,k\}$ iff $w\in W^{k}$, for each $k\in\mathbb{N}$, and $I_{w}(p)=1$ iff $w\in\bigcup_{i\in\mathbb{N}}W^{i}$. Then $\mathfrak{M}_{0}\models\varphi$ and so $\varphi\not\vdash_{L_{1}\otimes L_{2}}^{\ast}\bot$. On the other hand, it should be clear that $\mathfrak{M}\not\models\varphi$, for any finite model $\mathfrak{M}$. Hence, $L_{1}\otimes L_{2}$ does not have the global $d$-fmp. ∎
3.2 Local Consequence
For an $\mathcal{L}$-formula $\varphi$, consider the following set of subformulas:
$\Theta^{\mathbf{1}}(\varphi)=\{P(x)\mid P(x)\in\mathop{\textit{sub}}\varphi\}\cup\{\chi\in\mathop{\textit{sub}}\Box_{2}\psi\mid\Box_{2}\psi\in\mathop{\textit{sub}}\varphi\}.$
It can be thought of as all ‘atomic’ subformulas that have all the outermost $\Box_{1}$-operators removed. The set $\Theta^{\mathbf{2}}(\varphi)$ is defined symmetrically. For a set $\Phi$ of $\mathcal{L}$-formulas, let $\Theta^{\mathbf{i}}(\Phi)=\bigcup_{\varphi\in\Phi}\Theta^{\mathbf{i}}(\varphi)$.
We now define the notion of alternation depth of a formula using a converse operation to $\Theta^{\mathbf{i}}$. More precisely, for an $\mathcal{L}$-formula $\psi$ with surrogates for $\Box_{2}$-subformulas, we denote
$\psi\uparrow^{\mathbf{1}}=\psi[(\Box_{2}\chi)^{\mathbf{1}}/(\Box_{2}\chi)^{\mathbf{2}}\mid(\Box_{2}\chi)^{\mathbf{1}}\in\mathop{\textit{sub}}\psi];$
In other words, operation $\cdot\uparrow^{\mathbf{1}}$ restores the ‘meaning’ of the surrogates by re-introducing their $\Box_{2}$-modalities, but the replacement subformulas are built from surrogates for their own $\Box_{1}$-subformulas instead (and so contain no $\Box_{1}$-modalities themselves); $\psi\uparrow^{\mathbf{2}}$ is defined symmetrically. It should be clear that, if these operations are applied repeatedly, alternating between $\cdot\uparrow^{\mathbf{1}}$ and $\cdot\uparrow^{\mathbf{2}}$, then we arrive at the original formula without surrogates: more precisely, for any $\mathcal{L}$-formula $\varphi$, we have $\varphi=(\varphi^{\mathbf{1}})\uparrow^{\mathbf{1}}\uparrow^{\mathbf{2}}\cdots\uparrow^{\mathbf{i}}$, for some $\mathbf{i}\in\{\mathbf{1},\mathbf{2}\}$. Denote by $\textit{adp}^{\mathbf{1}}(\varphi)$ the minimum length of such a sequence; symmetrically, $\textit{adp}^{\mathbf{2}}(\varphi)$ is the smallest number of $\cdot\uparrow^{\mathbf{2}}$ and $\cdot\uparrow^{\mathbf{1}}$ operations such that $\varphi=(\varphi^{\mathbf{2}})\uparrow^{\mathbf{2}}\uparrow^{\mathbf{1}}\cdots\uparrow^{\mathbf{i}}$. For instance, consider $\varphi=\Box_{1}\Box_{2}p\ \land\ \Box_{2}\Box_{2}\Box_{1}r$. We then have the following:
$(\Box_{1}\Box_{2}p\ \land\ \Box_{2}\Box_{2}\Box_{1}r)^{\mathbf{1}}$ $\ \ =\ \ \Box_{1}(\Box_{2}p)^{\mathbf{1}}\ \land\ (\Box_{2}\Box_{2}\Box_{1}r)^{\mathbf{1}},$
$\bigl(\Box_{1}(\Box_{2}p)^{\mathbf{1}}\ \land\ (\Box_{2}\Box_{2}\Box_{1}r)^{\mathbf{1}}\bigr)\uparrow^{\mathbf{1}}$ $\ \ =\ \ \Box_{1}\Box_{2}p\ \land\ \Box_{2}\Box_{2}(\Box_{1}r)^{\mathbf{2}},$
$\bigl(\Box_{1}\Box_{2}p\ \land\ \Box_{2}\Box_{2}(\Box_{1}r)^{\mathbf{2}}\bigr)\uparrow^{\mathbf{2}}$ $\ \ =\ \ \Box_{1}\Box_{2}p\ \land\ \Box_{2}\Box_{2}\Box_{1}r.$
Thus, $\textit{adp}^{\mathbf{1}}(\varphi)=2$. Note that $\textit{adp}^{\mathbf{1}}{(\Box_{1}\Box_{2}p)}=1$ and $\textit{adp}^{\mathbf{1}}{(\Box_{2}\Box_{2}\Box_{1}r)}=2$. So, $\textit{adp}^{\mathbf{1}}{(\varphi)}$ is the maximal number of alternations between $\Box_{1}$ and $\Box_{2}$ in $\varphi$ starting from $\Box_{1}$. It should be clear that
$|\textit{adp}^{\mathbf{1}}{(\varphi)}-\textit{adp}^{\mathbf{2}}{(\varphi)}|\leq 1,$
depending on the leads of the longest alternation sequences in $\varphi$. Denote $\textit{adp}^{\mathbf{i}}(\Phi)=\textit{adp}^{\mathbf{i}}(\bigwedge\nolimits_{\varphi\in\Phi}\varphi)$. Let $\textit{adp}(\varphi)=\min(\textit{adp}^{\mathbf{1}}{(\varphi)},\textit{adp}^{\mathbf{2}}{(\varphi)})$ and $\textit{adp}(\Phi)=\textit{adp}(\bigwedge\nolimits_{\varphi\in\Phi}\varphi)$. Observe that, for any $\mathcal{L}$-formula,
$\text{if }\textit{adp}(\Phi)=\textit{adp}^{\mathbf{i}}(\Phi),\text{ then }\textit{adp}(\Theta^{\mathbf{i}}(\Phi))=\max\{0,\textit{adp}(\Phi)-1\}\text{ and }\textit{adp}(\Theta^{\mathbf{i}}(\Phi))=\textit{adp}^{\mathbf{3-i}}(\Theta^{\mathbf{i}}(\Phi)).$ ($\dagger$)
In addition to adp, we define the modal $1$-depth $\textit{md}^{\mathbf{1}}(\varphi)$ of $\varphi$ as the maximum number of nested $\Box_{1}$ (ignoring $\Box_{2})$; the modal $2$-depth $\textit{md}^{\mathbf{2}}(\varphi)$ is defined symmetrically. We also set $\textit{md}^{\mathbf{i}}(\Phi)=\textit{md}^{\mathbf{i}}(\bigwedge\nolimits_{\varphi\in\Phi}\varphi)$.
Finally, for an $\mathcal{L}$-formula $\varphi$, consider
$\boldsymbol{Q}_{i}(\varphi)=\bigl\{\boldsymbol{q}\text{ a $\Theta^{\mathbf{i}}(\varphi)$-quasistate such that }\not\vdash_{L_{1}\otimes L_{2}}\neg\hat{\boldsymbol{q}}\bigr\}$,
and, as before, denote by $\hat{\boldsymbol{Q}}_{i}(\varphi)$ the disjunction of all $\hat{\boldsymbol{q}}$, for $\boldsymbol{q}\in\boldsymbol{Q}_{i}(\varphi)$. With these definitions at hand, we can formulate our criterion for deciding local consequence, as claimed in Theorem 3.1.
Lemma 3.6.
For all $\mathcal{L}$-formulas $\varphi$ and $i=1,2$ with $\textit{adp}(\varphi)=\textit{adp}^{\mathbf{i}}{(\varphi)}$, the following are equivalent:
$\not\vdash_{L_{1}\otimes L_{2}}\varphi$;
there is a $d$-model $\mathfrak{M}$ based on a $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$-frame such that $\mathfrak{M}\not\models\varphi$;
$\not\vdash_{L_{i}}\displaystyle\Box_{i}^{\leq\textit{md}^{\mathbf{i}}(\varphi)}(\hat{\boldsymbol{Q}}_{i}(\varphi))^{\mathbf{i}}\to\varphi^{\mathbf{i}}$.
Completeness of $L_{1}\otimes L_{2}$ for $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$ claimed in Theorem 3.1 then follows immediately. For decidability, observe that, while the definition of $\boldsymbol{Q}_{i}(\varphi)$ refers to $\vdash_{L_{1}\otimes L_{2}}$ and thus may appear circular, by property ($\dagger$3.2), the $\Theta^{\mathbf{i}}(\varphi)$-formulas have smaller alternation depth, and so we can apply the criterion in Lemma 3.6 recursively to the simpler formulas to decide whether the respective $\boldsymbol{q}$ should be included in the set $\boldsymbol{Q}_{i}(\varphi)$. This recursion terminates as we eventually reach formulas of alternation depth 0, which, by definition, belong to either $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$, where, by assumption, the consequence relation is decidable.
We prove Lemma 3.6 for expanding domains (the case of constant domains is similar). We require some additional definitions.
Definition 3.7.
Let $\Phi$ be a set of $\mathcal{L}$-formulas closed under subformulas, For $i=1,2$ and $h\leq\textit{md}^{\mathbf{i}}(\Phi)$, denote $\Gamma_{h}^{\mathbf{i}}(\Phi)=\{\psi\in\Phi\mid\textit{md}^{\mathbf{i}}(\psi)\leq h\}$. Since $\Gamma^{\mathbf{i}}_{h}(\Phi)=\emptyset$ for $h<0$, there is only one such $\Gamma^{\mathbf{i}}_{h}(\Phi)$-type, $\emptyset$. For a type $t$, let $t|_{\Phi}$ be the restriction of $t$ to $\Phi$, that is, $t\cap\{\psi,\neg\psi\mid\psi\in\Phi\}$, and denote $\boldsymbol{q}|_{\Phi}=\{t|_{\Phi}\mid t\in\boldsymbol{q}\}$.
Let $(W,R_{i})\in\mathcal{C}_{i}$. For $w,w^{\prime}\in W$, let $\textit{dist}_{R_{i}}(w,w^{\prime})$ be the length of the shortest $R_{i}$-path from $w$ to $w^{\prime}$. A tapered $\Phi$-run through $((W,R_{i}),w_{0})$ is a function $r$ from an $R_{i}$-closed subset of $W$, denoted by $\mathop{\mathsf{dom}}r$, such that each $r(w)$ is a $\Gamma_{h(w)}^{\mathbf{i}}(\Phi)$-type, where $h(w)=\textit{md}^{\mathbf{i}}(\Phi)-\textit{dist}_{R_{i}}(w_{0},w)$. For $w\in\mathop{\mathsf{dom}}r$, run $r$ is
  • $i$-coherent at $w$ if $\Box_{i}\psi\in r(w)$ implies $\psi\in r(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$;
  • $i$-saturated at $w$ if $\neg\Box_{i}\psi\notin r(w)$ whenever $\psi\in r(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$. [^3] We use $\neg\Box_{i}\psi\notin r(w)$ as a convenient replacement for either $\Box_{i}\psi\notin\Gamma_{h(w)}^{\mathbf{i}}(\Phi)$ or $\Box_{i}\psi\in r(w)$.
We say that $r$ is $i$-coherent on $W^{\prime}\subseteq W$ if $r$ is $i$-coherent at each $w\in\mathop{\mathsf{dom}}r\cap W^{\prime}$ and $i$-coherent if $r$ is $i$-coherent on $W$; similarly for $r$ being $i$-saturated.
A tapered $i$-quasimodel $\mathfrak{Q}$ for $\Phi$ is a tuple of the form $((W,R_{i}),w_{0},\mathfrak{R})$, where $(W,R_{i})\in\mathcal{C}_{i}$, $w_{0}\in W$, and $\mathfrak{R}$ is a set of tapered $\Phi$-runs through $((W,R_{i}),w_{0})$ such that each $r\in\mathfrak{R}$ is $i$-coherent and $i$-saturated and each $\mathfrak{R}(w)$ is a $\Phi$-quasistate, for $w\in W$, with $\mathfrak{R}_{w}=\{r\in\mathfrak{R}\mid w\in\mathop{\mathsf{dom}}r\}$ and $\mathfrak{R}(w)=\{r(w)\mid r\in\mathfrak{R}_{w}\}$.
Definition 3.8.
We inductively define a tapered $1$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$, where $(W,R_{1})\in\mathcal{C}_{1}$ and $(W\setminus\{w_{0}\},R_{2})\in\mathcal{C}_{2}$ with a distinguished world $w_{0}\in W$, which is not in the domain or range of $R_{2}$, and $\mathfrak{R}$ is a set of tapered runs that are $1$- and $2$-coherent as well as $1$-saturated on $W$ but $2$-saturated only on $W\setminus\{w_{0}\}$. A tapered $2$-cactus is defined symmetrically. See Fig. 1, where we show the structure of the frame.
[…]
degenerate
2-cacti
degenerate
1-cacti
$\textit{adp}(\varphi^{w_{3}})=1$
$\textit{md}^{\mathbf{1}}(\varphi^{w_{3}})=1$
$\Phi^{w_{3}}=\$
$\Theta^{\mathbf{2}}(\Phi^{w_{0}})\cap\Gamma^{\mathbf{2}}_{0}(\Phi^{w_{0}})=\$
$\{p,\Box_{1}\neg r,\dots\}$
Figure 1: A fully grafted tapered $1$-cactus satisfying $\varphi=\Diamond_{1}\Diamond_{2}p\land\Diamond_{2}\Diamond_{2}\Diamond_{1}r$: solid arrows with dark background for $R_{1}$-components and dotted arrows with light background for $R_{2}$-components.
basis: $\Phi=\emptyset$.
A tapered ‘degenerate’ $1$-cactus $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ for $\emptyset$ is constructed by first taking sufficiently large disjoint unions $(W,R_{1})$ and $(W^{\prime},R_{2}^{\prime})$ of copies of $\mathfrak{F}_{1}\in\mathcal{C}_{1}$ and $\mathfrak{F}_{2}\in\mathcal{C}_{2}$, respectively, such that $|W|-1=|W^{\prime}|$. We then take any $w_{0}\in W$ and obtain the required $(W,R_{1},R_{2})$ by applying to $R_{2}^{\prime}$ the bijection between $W\setminus\{w_{0}\}$ and $W^{\prime}$. Finally, $\mathfrak{R}$ consists of a single tapered $\emptyset$-run $r$ with $r(w)=\emptyset$, for all $w\in W$, which is trivially $1$- and $2$-coherent as well as $1$- and $2$-saturated.
induction step: $\Phi\neq\emptyset$.
Let $\mathfrak{Q}=((W^{0},R_{1}^{0}),w_{0},\mathfrak{R}^{0})$ be a tapered $1$-quasimodel for $\Phi$. For each $v\in W^{0}$, denote
$\Phi^{v}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(v)}(\Phi),\quad\text{ where }h(v)=\textit{md}^{\mathbf{1}}(\Phi)-\textit{dist}_{R_{1}^{0}}(w_{0},v).$
For each $v\in W^{0}\setminus\{w_{0}\}$, take a tapered $2$-cactus $\mathfrak{C}^{v}=((W^{v},R_{1}^{v},R_{2}^{v}),v,\mathfrak{R}^{v})$ for $\Phi^{v}$ that is compatible with $\mathfrak{Q}$ in the sense that
$W^{0}\cap W^{v}=\{v\}\quad\text{ and }\quad\mathfrak{R}^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}=\mathfrak{R}^{v}(v).$
We also assume that the $W^{v}$ are disjoint. Then $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ is a tapered $1$-cactus $\mathfrak{C}$ for $\Phi$ if it is obtained by grafting the $\mathfrak{C}^{v}$, for $v\in W^{0}\setminus\{w_{0}\}$, onto $\mathfrak{Q}$. More precisely, we say that $((W^{0},R_{1}^{0},\emptyset),w_{0},\mathfrak{R}^{0})$ is an initial stock and $W^{0}$ its thorns. Then, given a stock $((W,R_{1},R_{2}),w_{0},\mathfrak{R})$ and a thorn $v\in W$, the grafting of $\mathfrak{C}^{v}$ onto $v$ is a stock $((W\cup W^{v},R_{1}\cup R_{1}^{v},R_{2}\cup R_{2}^{v}),w_{0},\mathfrak{R}^{\prime})$, where
$\mathfrak{R}^{\prime}\ \ =\ \ \mathfrak{R}\setminus\mathfrak{R}_{v}\ \ \cup\ \ \bigl\{r\cup r^{v}\mid r\in\mathfrak{R}_{v}\text{ and }r^{v}\in(\mathfrak{R}^{v})_{v}\text{ with }r(v)|_{\Theta^{\mathbf{1}}(\Phi)}=r^{v}(v)\bigr\}\ \ \cup\ \ \mathfrak{R}^{v}\setminus(\mathfrak{R}^{v})_{v}.$
Observe that since $v$ in $\mathfrak{C}^{v}$ is not $R_{1}$-connected to any world, the $R_{1}$-components in $\mathfrak{Q}$ and $\mathfrak{C}^{v}$ are disjoint, and so the frame in the resulting stock belongs to $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$. It then follows that the runs are $1$-coherent, 11-saturated and $2$-coherent on $W\cup W^{v}$. The runs are also $2$-saturated on $W^{\prime}\cup W^{v}$ if they are $2$-saturated on $W^{\prime}\subseteq W$ in the given stock.
A fully grafted tapered $i$-cactus $((W,R_{1},R_{2}),\mathfrak{R})$ for $\Phi$ results from grafting a compatible tapered (3i)(3-i)-cactus $\mathfrak{C}^{w_{0}}$ for $\Phi^{w_{0}}=\Theta^{\mathbf{i}}(\Phi)$ onto the world $w_{0}$ of a tapered $i$-cactus $((W^{0},R_{1}^{0},R_{2}^{0}),w_{0},\mathfrak{R}^{0})$ for $\Phi$. Runs in a fully grafted tapered cactus are $i$-coherent and $i$-saturated for both $i=1,2$ on the whole $W$.
By induction on alternation depth, we show the following claim, which implies Lemma 3.6.
Claim 3.9.
For all $\mathcal{L}$-formulas $\varphi$ and $i=1,2$ with $\textit{adp}(\varphi)=\textit{adp}^{\mathbf{i}}{(\varphi)}$, each of (L1)(L3) holds iff
there is a (fully grafted) tapered $i$-cactus $((W,R_{1},R_{2}),\mathfrak{R})$ for $\mathop{\textit{sub}}\varphi$ with $\varphi\notin r(w_{0})$, for some $r\in\mathfrak{R}$.
Proof.
The basis of induction, $\textit{adp}(\varphi)=0$, is trivial as the formula belongs to $\mathcal{L}_{1}$ or $\mathcal{L}_{2}$ (or both, if it has no modalities). For the inductive step, consider an $\mathcal{L}$-formula $\varphi$ and suppose that $\textit{adp}(\varphi)=k>0$. Assume that $\textit{adp}^{\mathbf{1}}(\varphi)=k$ (the other case is symmetrical) and that the claim has been shown for all formulas $\psi$ with $\textit{adp}^{\mathbf{2}}(\psi)<k$, which, by ($\dagger$3.2), includes all $\psi\in\Theta^{\mathbf{1}}(\varphi)$.
$\textbf{(L2)}\Rightarrow\textbf{(L1)}$ is by soundness of $\vdash_{L_{1}\otimes L_{2}}$, and $\textbf{(L1)}\Rightarrow\textbf{(L3)}$ because $\vdash_{L_{i}}$ is closed under substitution.
$\textbf{(L3)}\Rightarrow\textbf{(L4)}$ By Kripke completeness of $L_{1}$, there is an xd-model $\mathfrak{M}^{0}$ based on $(W^{0},R^{0}_{1})\in\mathcal{C}_{1}$ such that $\mathfrak{M}^{0},w_{0}\models\Box_{1}^{\smash{\leq\textit{md}^{\mathbf{1}}(\varphi)}}(\hat{\boldsymbol{Q}}_{1}(\varphi))^{\mathbf{1}}$ and $\mathfrak{M}^{0},w_{0}\not\models\varphi^{\mathbf{1}}$ for some $w_{0}\in W^{0}$. Let $\Phi=\mathop{\textit{sub}}\varphi$. For each $v\in W^{0}$, let $h(v)=\textit{md}^{\mathbf{1}}(\Phi)-\textit{dist}_{R^{0}_{1}}(w_{0},v)$. Then, for each $v\in W^{0}$ and $e\in\Delta^{0}_{v}$, we define the $\Gamma^{\mathbf{1}}_{h(v)}(\Phi)$-type $t_{v,e}$ of $e$ at $v$ such that
$\mathfrak{M},v\models t_{v,e}^{\mathbf{1}}[e].$
Next, we construct a set $\mathfrak{R}^{0}$ of tapered $\Phi$-runs on $(W^{0},R^{0}_{1})$: for each $e\in\bigcup_{v\in W^{0}}\Delta^{0}_{v}$, we consider $r_{e}$ that maps each $v\in W^{0}$ with $e\in\Delta^{0}_{v}$ to $t_{v,e}$. It follows that $\mathfrak{Q}=((W^{0},R^{0}_{1}),w_{0},\mathfrak{R}^{0})$ is a tapered 1-quasimodel for $\Phi$ such that $\varphi\notin t_{w_{0},e}$, for some $e\in\Delta_{w_{0}}^{0}$.
For each $v\in W^{0}$ with $h(v)\geq 0$, we consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{v}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{v})^{\mathbf{1}}$. Observe that $\boldsymbol{q}^{v}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(v)$. Let $\Phi^{v}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(v)}(\Phi)$ and consider $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}$, which is a $\Phi^{v}$-quasistate with $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}$. Let $\varphi^{v}$ be its realisability sentence. Since $\textit{adp}^{\mathbf{2}}(\varphi^{v})<k$, we apply the induction hypothesis to the $\mathcal{L}$-formula $\varphi^{v}$ and obtain a tapered 2-cactus $\mathfrak{C}^{v}=((W^{v},R_{1}^{v},R_{2}^{v}),v,\mathfrak{R}^{v})$ for $\Phi^{v}$ (strictly speaking, for $\mathop{\textit{sub}}\varphi^{v}$ but the only difference is that $\mathop{\textit{sub}}\varphi^{v}$ contains Boolean combinations of quantified formulas from $\Phi^{v}$) such that $\boldsymbol{q}^{v}|_{\smash{\Gamma^{\mathbf{1}}_{h(v)}(\Phi)}}=\mathfrak{R}^{v}(v)$. For each $v\in W^{0}$ with $h(v)<0$, we take a tapered 2-cactus $\mathfrak{C}^{v}$ for $\emptyset$. Next, for each $v\in W^{0}\setminus\{w_{0}\}$, we graft the tapered 2-cactus $\mathfrak{C}^{v}$ onto $v$ in the 1-quasimodel $\mathfrak{Q}$ to obtain a tapered 1-cactus $\mathfrak{C}$. A fully grafted tapered 1-cactus can be obtained by grafting, in addition, $\mathfrak{C}^{w_{0}}$ onto $w_{0}$ in $\mathfrak{C}$.
$\textbf{(L4)}\Rightarrow\textbf{(L2)}$ Let $\mathfrak{C}=((W,R_{1},R_{2}),\mathfrak{R})$ be a fully grafted tapered $i$-cactus for $\mathop{\textit{sub}}\varphi$ with $\varphi\notin r(w_{0})$, for some $r\in\mathfrak{R}$. We construct $\mathfrak{M}$ based on a $\mathcal{C}_{1}\otimes\mathcal{C}_{2}$-frame $(W,R_{1},R_{2})$ and domains $\Delta_{w}=\mathfrak{R}_{w}$, for $w\in W$: we set $\mathfrak{M},w\models P[r]$ iff $P(x)\in r(w)$, for each unary predicate $P$, $w\in W$ and $r\in\mathfrak{R}_{w}$. By induction on the structure of formulas, we show that, for each $w\in W$, $r\in\mathfrak{R}_{w}$, $\chi\in\{\psi,\neg\psi\mid\psi\in r(w)\}$ and $i=1,2$,
$\mathfrak{M},w\models\chi[r]\qquad\text{ iff }\qquad\chi^{\mathbf{i}}\in(r(w))^{\mathbf{i}}.$
The basis of induction is by definition. The cases of the Boolean connectives and quantifiers are immediate from the definition of types and quasistates. The cases of the modalities follow from the fact that the runs in $\mathfrak{C}$ are $i$-coherent and $i$-saturated. Thus, $\mathfrak{M},w_{0}\not\models\varphi$. ∎
Theorem 3.10.
For $d\in\{\textit{xd},\textit{cd}\}$, if $L_{1}$ and $L_{2}$ are one-variable modal logics with the $d$-fmp, then $L_{1}\otimes L_{2}$ also has the $d$-fmp.
The proof of this result follows from the proof of Claim 3.9: if $L_{1}$ and $L_{2}$ have the finite model property, then the tapered cacti (including the degenerate cacti for the basis) constructed for $\textbf{(L3)}\Rightarrow\textbf{(L4)}$ are all finite, both in the number of worlds and the number of runs; but then such a finite tapered cactus results in a finite model constructed in $\textbf{(L4)}\Rightarrow\textbf{(L2)}$ to satisfy an $\mathcal{L}$-formula.
4 Fusions of One-Variable Logics with Equality
We show that once equality is allowed in one‑variable modal logics, the transfer results established in Section 3 no longer hold. Remarkably, the counterexamples we introduce next arise quite naturally. The propositional modal logic $\mathbf{Diff}$ of the ‘elsewhere’ operator was introduced by Von Wright [32] and is determined by the class $\mathcal{D}$ of all disjoint unions of frames $(W,R)$ with $R=\{(w,v)\in W^{2}\mid w\not=v\}$; see also [8]. We are interested here in the one-variable modal logics $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ and $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ with $\mathcal{D}_{\text{fin}}$ the class of disjoint unions of finite frames in $\mathcal{D}$. Note that the propositional logic Diff has the ffmp, and so is complete for both $\mathcal{D}$ and $\mathcal{D}_{\text{fin}}$. On the other hand, while one can easily prove the fmp for $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$, the logic $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ does not have the ffmp, and so $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}\subsetneq\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$ [20]. It has been observed [19, 20] that both logics can be reduced in polytime to the two-variable fragment of first-order logic with counting on arbitrary and finite domains, respectively, and so they are decidable in coNExpTime.
Theorem 4.1.
Both $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ and $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\textup{fin}}$ are decidable but $\mathop{\textup{Log}_{\textit{cd}}^{=}}(\mathcal{D}\otimes\mathcal{D}_{\textup{fin}})$ is undecidable.
Proof.
The proof is by reduction of the problem of solving Diophantine equations. For arbitrary polynomials $g,h$ with coefficients from $\mathbb{N}\setminus\{0\}$, it is undecidable whether $g=h$ has a solution in $\mathbb{N}\setminus\{0\}$ [7]. Each polynomial equation can be rewritten into a set of elementary equations
$y=n,\qquad y=z_{1}+z_{2},\qquad y=z_{1}\times z_{2}$
with variables $y,z_{1},z_{2}$ and $n\in\mathbb{N}\setminus\{0\}$. So we assume that a set of such equations $E=\{e_{1},\ldots,e_{n}\}$ is given. We construct a formula $\varphi_{E}$ which is satisfiable iff $E$ has a solution in $\mathbb{N}\setminus\{0\}$. Let the modal operators $\Box$ and $\Box_{C}$ be interpreted over frames with accessibility relations $R$ in $\mathcal{D}$ and $R_{C}$ in $\mathcal{D}_{\text{fin}}$, respectively. We set $\Box^{+}\varphi=\varphi\wedge\Box\varphi$, and likewise for $\Box_{C}$. Formula $\varphi_{E}$ will hardly need any alternation between modalities, the maximal alternation being of the form $\Box\Box_{C}\chi$. In fact, the only contribution of $\mathcal{D}_{\text{fin}}$ is to supply the required finite cardinality of the connected components of its frames.
When discussing the construction of $\varphi_{E}$, we assume a world $w$ in the set $W$ of worlds reachable along $R$. To $W$ attached are the models for $\Box_{C}$ with the $R_{C}$-connected component of $w$ denoted by $W_{w}$. For a unary predicate $P$ and $w\in W$, we denote by $|P|_{w}$ the cardinality of the extension of $P$ at $w$. In the sequel, for a closed formula $I$ and a unary predicate $P$, we use formula $\textsf{card}(I,P)$ defined as follows:
$\Box^{+}\bigl[I\ \ \to\ \ \Box_{C}^{+}\forall x\,\bigl((x=c)\rightarrow\Box_{C}(x\neq c)\bigr)\ \ \land\ \ \forall x\,\bigl(P(x)\leftrightarrow\Diamond_{C}^{+}(x=c)\bigr)\bigr],$
where $c$ is a fresh constant. Formula $\textsf{card}(I,P)$ ensures that, at any world $w$ where $I$ holds, the number of worlds in the $R_{C}$-connected component of $w$ coincides with the number of domain elements in the extension of $P$ at $w$, that is, $|W_{w}|=|P|_{w}$; see Fig. 2 a), where horizontal lines represent the domain.
For each variable $y$ used on the left- or right-hand side of equations in $E$, we take unary predicates $P_{y}$ and $C_{y}$. Intuitively, $|P_{y}|_{w}$ encodes the value of $y$, while $C_{y}$ is used to ensure that the value of $y$ is finite. To this end, we first mark a single world $w\in W$ with $I_{y}=\exists x\,C_{y}(x)$:
$\Diamond^{+}(I_{y}\wedge\Box\neg I_{y}).$
Next, at the world $w\in W$ marked with $I_{y}$, we use formula $\textsf{card}(I_{y},P_{y})$ to ensure that $|P_{y}|_{w}=|W_{w}|$. It follows, in particular, that $|P_{y}|_{w}$ is finite. Finally, the following conjuncts guarantee that each $P_{y}$ is non-empty (only solutions in $\mathbb{N}\setminus\{0\}$ are admitted) and constant on $W$:
[…]
If $e\in E$ is of the form $y=n$, then we take $n$ constants $c_{1}^{e},\ldots,c_{n}^{e}$ interpreted differently and make the set of their interpretations the extension of $P_{y}$:
$\forall x\,\bigl(P_{y}(x)\leftrightarrow\bigvee_{1\leq i\leq n}(x=c_{i}^{e})\bigr)\land\bigwedge_{1\leq i<j\leq n}\hskip-5.0pt(c_{i}^{e}\neq c_{j}^{e}).$
If $e$ is of the form $y=z_{1}+z_{2}$, then we introduce two unary predicates $Q^{e}_{1}$ and $Q^{e}_{2}$ and, for both $i=1,2$, add formula $\textsf{card}(I_{z_{i}},Q^{e}_{i})$, which ensures $|P_{z_{i}}|_{w_{i}}=|Q^{e}_{i}|_{w_{i}}$ at the world $w_{i}\in W$ marked with $I_{z_{i}}$. Next, we state that $Q^{e}_{1}$ and $Q^{e}_{2}$ are disjoint at all worlds $w\in W$ and empty at worlds $w\in W$ where the respective $I_{z_{i}}$ does not hold:
$\Box^{+}\forall x\,\bigl(Q^{e}_{1}(x)\to\Box^{+}\neg Q^{e}_{2}(x)\bigr),\qquad\Box^{+}\bigl(\exists x\,Q^{e}_{1}(x)\to I_{z_{1}}\bigr)\quad\text{ and }\quad\Box^{+}\bigl(\exists x\,Q^{e}_{2}(x)\to I_{z_{2}}\bigr).$
Finally, we add the conjunct that ensures that $P_{y}$ is the union of $Q^{e}_{1}$ and $Q^{e}_{2}$:
$\forall x\,\bigl(P_{y}(x)\leftrightarrow\Diamond^{+}Q^{e}_{1}(x)\vee\Diamond^{+}Q^{e}_{2}(x)\bigr),$
which implies that equation $e$ holds in the sense that $|P_{y}|_{w}=|P_{z_{1}}|_{w}+|P_{z_{2}}|_{w}$, for all worlds $w\in W$.
[…]
Figure 2: a) Formula $\textsf{card}(I,P)$. b) Equation $e$ of the form $y=z_{1}\times z_{2}$ with $z_{1}=3$ and $z_{2}=2$.
If $e$ is of the form $y=z_{1}\times z_{2}$, then we introduce a unary predicate $C^{e}$ and a constant $c^{e}$ and let $J^{e}(x)=C^{e}(x)\land(c^{e}=x)$. First, we ensure $|P_{z_{2}}|_{w}$ is equal to the number of worlds in $W$ satisfying $\exists x\,J^{e}(x)$:
[…]
where, due to the second conjunct, $J^{e}(x)$ cannot hold on the same domain element more than once in $W$. Next, we introduce a unary predicate $Q^{e}$ and add formulas $\textsf{card}(\exists x\,J^{e}(x),P_{z_{1}})$ and $\textsf{card}(\exists x\,J^{e}(x),Q^{e})$ to ensure that $|P_{z_{1}}|_{w}=|Q^{e}|_{w}$ for all worlds $w\in W$ with $\exists x\,J^{e}(x)$. Then we state that the $Q^{e}$ are mutually disjoint at different worlds $w\in W$ and empty at worlds $w\in W$ where $\exists x\,J^{e}(x)$ does not hold:
[…]
So now we have $|P_{z_{1}}|_{w}$ (encoded as $|Q^{e}|_{w}$) in exactly $|P_{z_{2}}|_{w}$-many worlds and take the sum by adding the conjunct
$\forall x\,\bigl(P_{y}(x)\leftrightarrow\Diamond^{+}Q^{e}(x)\bigr),$
which implies that equation $e$ holds in the sense that $|P_{y}|_{w}=|P_{z_{1}}|_{w}\times|P_{z_{2}}|_{w}$, for all $w\in W$; see Fig. 2 b), where only worlds in $W$ and their domains are shown.
This completes the reduction. ∎
Corollary 4.2.
For $d\in\{xd,cd\}$, decidability, decidability of global consequence, and recursive axiomatisability are not preserved under the mapping
$(\mathop{\textup{Log}^{=}_{d}}\mathcal{C}_{1},\ \mathop{\textup{Log}^{=}_{d}}\mathcal{C}_{2})\mapsto\mathop{\textup{Log}^{=}_{d}}(\mathcal{C}_{1}\otimes\mathcal{C}_{2}),$
for classes $\mathcal{C}_{1},\mathcal{C}_{2}$ of frames closed under disjoint unions.
Proof.
Theorem 4.1 does not depend on the constant domain semantics since $\mathcal{D}$ and $\mathcal{D}_{\text{fin}}$ do not feel the difference between expanding and constant domains. Likewise, the universal modality is definable in $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}$ and $\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\text{fin}}$, and so decidability of global consequence is not preserved either. Recursive axiomatisability is not preserved since we have shown that $\mathop{\textup{Log}_{\textit{cd}}^{=}}(\mathcal{D}\otimes\mathcal{D}_{\text{fin}})$ is not recursively enumerable. ∎
We next use a reduction of Minsky machines introduced for undecidability proofs in temporal first-order logic [9] to show the following result for the global consequence.
Theorem 4.3.
Let $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ be classes of Kripke frames that are closed under disjoint unions and contain non-trivial frames. Then the global consequence for $\mathop{\textup{Log}_{\textit{cd}}^{=}}(\mathcal{C}_{1}\otimes\mathcal{C}_{2})$ is undecidable.
A (two-counter) Minsky machine [26] is a pair $M=(Q,I)$, where $Q=\{q_{0},\ldots,q_{N}\}$ is a set of states and $I=\{I_{0},\ldots,I_{N\mbox{-}1}\}$ is a set of instructions. We assume that $q_{0}$ is the initial state, that $q_{N}$ is the halting state, and that, whenever $M$ is in state $q_{i}$, with $0\leq i<N$, the instruction to be executed is $I_{i}$. A machine has two registers, $r_{1}$ and $r_{2}$; at every instance of a computation, the register $r_{k}$, with $k\in\{1,2\}$, holds a non-negative integer $v_{k}$, called the value of $r_{k}$. Instructions in $I$ are of two types:
  • $(r_{k},q_{j})$ – add $1$ to $v_{k}$ and go to state $q_{j}$;
  • $(r_{k},q_{j},q_{l})$ – if $v_{k}>0$, subtract $1$ from $v_{k}$ and go to $q_{j}$; else, leave $v_{k}$ unchanged and go to $q_{l}$.
A configuration of $M$ is a triple $(q,v_{1},v_{2})$, where $q\in Q$ and $v_{1}$, $v_{2}$ are values of the registers. If $(q_{i},v_{1},v_{2})$ and $(q_{j},v^{\prime}_{1},v^{\prime}_{2})$ are configurations, we write $(q_{i},v_{1},v_{2})\Rightarrow_{M}(q_{j},v^{\prime}_{1},v^{\prime}_{2})$ if one of the following applies:
  • $I_{i}=+(r_{k},q_{j})$, $v^{\prime}_{k}=v_{k}+1$, and $v^{\prime}_{3-k}=v^{\phantom{}}_{3-k}$;
  • $I_{i}=-(r_{k},q_{j},q_{l})$, $v_{k}>0$, $v^{\prime}_{k}=v_{k}-1$, and $v^{\prime}_{3-k}=v^{\phantom{}}_{3-k}$;
  • $I_{i}=-(r_{k},q_{l},q_{j})$, $v^{\prime}_{k}=v_{k}=0$, and $v^{\prime}_{3-k}=v^{\phantom{}}_{3-k}$.
An input to $M$ is a pair $(n,m)\in\mathbb{N}\times\mathbb{N}$. The computation of $M$ on input $(n,m)$ is the (unique) longest sequence of configurations $(q^{0},v^{0}_{1},v^{0}_{2})\Rightarrow_{M}(q^{1},v^{1}_{1},v^{1}_{2})\Rightarrow_{M}\ldots$, where $q^{0}=q_{0}$, $v^{0}_{1}=n$, and $v^{0}_{2}=m$.
We reduce the undecidable halting problem for Minsky machines (decide, given a Minsky machine, whether it halts on $(0,0)$) to our problem. To that end, we use a language containing modalities $\Box_{1}$ and $\Box_{2}$, proposition letters $q_{1},\ldots,q_{N}$ for the states of a machine, unary predicate letters $V_{1}$ and $V_{2}$ for the values of the registers, proposition letter $p$ for making sure that we always transition to a distinct world, and a (non-rigid) constant symbol $c$.
We use the modal operators $\Box\Box$ and $\Diamond\Diamond$ defined in the proof of Theorem 3.5. For convenience, we define the following abbreviations:
$\textsf{counter}(\psi,\psi^{\prime})$ $=\forall x\,\bigl(\psi(x)\to\mbox{$\Box$}\hskip-7.7778pt\hskip 1.00006pt\mbox{$\Box$}\psi^{\prime}(x)\bigr)\wedge\forall x\,\bigl(\mbox{$\Diamond$}\hskip-6.66669pt\hskip 1.00006pt\mbox{$\Diamond$}\psi^{\prime}(x)\to\psi(x)\bigr),$ $\text{ for open formulas $\psi(x)$ and $\psi^{\prime}(x)$},$
[…]
We first describe the initial configuration of $M$ on $(0,0)$:
$\varphi_{\mathit{init}}$ $\ \ =\ \ q_{0}\wedge\neg\exists x\,V_{1}(x)\wedge\neg\exists x\,V_{2}(x).$
Next, we say that, at every step of the computation, $M$ is in at most one, non-halting, state:
$\varphi_{Q}$ $\ \ =\ \ \bigwedge_{1\leq i<j<N}\neg(q_{i}\land q_{j})\quad\land\quad\neg q_{N}.$
The effect of $M$’s instructions is described by $\varphi_{I}$, the conjunction of $\varphi_{i}$, for $1\leq i<N$, defined as follows:
  • if $I_{i}=+(r_{k},q_{j})$, then define $\varphi_{i}$ to be the formula
    $q_{i}\wedge\neg V_{k}(c)\ \to\ \textsf{next}(q_{j})\ \wedge\ \textsf{counter}(V_{k}(x)\vee(x=c),V_{k}(x))\ \wedge\ \textsf{counter}(V_{3-k}(x),V_{3-k}(x));$
  • if $I_{i}=-(r_{k},q_{j},q_{l})$, then define $\varphi_{i}$ to be the conjunction of formulas
    $q_{i}\wedge V_{k}(c)$ $\ \to\ \textsf{next}(q_{j})\ \wedge\ \textsf{counter}(V_{k}(x)\wedge(x\neq c),V_{k}(x))\ \wedge\ \textsf{counter}(V_{3-k}(x),V_{3-k}(x)),$
    $q_{i}\wedge\neg\exists x\,V_{k}(x)$ $\ \to\ \textsf{next}(q_{l})\ \wedge\ \textsf{counter}(V_{k}(x),V_{k}(x))\ \wedge\ \textsf{counter}(V_{3-k}(x),V_{3-k}(x)).$
Lemma 4.4.
Let $\mathscr{C}_{1}$ and $\mathscr{C}_{2}$ be classes of Kripke frames that are closed under disjoint unions and contain non-trivial frames. Then a Minsky machine $M$ does not halt on $(0,0)$ iff there is a cd-model $\mathfrak{M}$ over a frame from $\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ and $\mathfrak{M},w_{0}\models\varphi_{\mathit{init}}$, for some $w_{0}$ in $\mathfrak{M}$.
Proof.
$(\Rightarrow)$ Let
$\begin{array}[]{lcl}C_{0}&=&(q^{0},0,0)\Rightarrow_{M}(q^{1},v^{1}_{1},v^{1}_{2})\Rightarrow_{M}\ldots\Rightarrow_{M}(q^{n},v^{n}_{1},v^{n}_{2})\Rightarrow_{M}\ldots\end{array}$
be the non-terminating computation of $M$ on $(0,0)$. We define a cactus-like model $\mathfrak{M}$ based on a frame $\mathfrak{F}\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ with a constant domain $\mathbb{N}$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ and $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$, for some world $w^{0}$.
Let $\mathfrak{G}_{1}=(U_{1},S_{1})\in\mathscr{C}_{1}$ and $\mathfrak{G}_{2}=(U_{2},S_{2})\in\mathscr{C}_{2}$ be non-trivial frames; thus, $\mathfrak{G}_{1}$ contains worlds $w$ and $v$ such that $w\neq v$ and $wS_{1}v$; similarly, $\mathfrak{G}_{2}$ contains worlds $s$ and $t$ such $s\neq t$ and $sS_{2}t$. The frame $\mathfrak{F}=(W,R_{1},R_{2})\in\mathscr{C}_{1}\otimes\mathscr{C}_{2}$ is a limit cactus constructed from a sequence $\{\mathfrak{G}_{1}^{i}=(U_{1}^{i},S_{1}^{i})\mid i\in\mathbb{N}\}$ of disjoint copies of $\mathfrak{G}_{1}$ and a sequence $\{\mathfrak{G}_{2}^{i}=(U_{2}^{i},S_{2}^{i})\mid i\in\mathbb{N}\}$ of disjoint copies of $\mathfrak{G}_{2}$ (starting with $\mathfrak{G}^{0}_{1}$). To simulate the computation $C_{0}$, we shall use infinite sequences of worlds from $\mathfrak{F}$ of the form
$\begin{array}[]{c}w^{0}R_{1}v^{0}(=s^{0})R_{2}t^{0}(=w^{1})R_{1}v^{1}(=s^{1})R_{2}\ldots\end{array}$
where $w^{i},v^{i}\in U^{i}_{1}$ and $s^{i},t^{i}\in U^{i}_{2}$, and also $w^{i}\neq v^{i}$ and $s^{i}\neq t^{i}$, for all $i\in\mathbb{N}$ (by the assumption about non-triviality of $\mathfrak{G}_{1}$ and $\mathfrak{G}_{2}$, the set of such sequences is non-empty). The $n$th configuration of $C_{0}$ will be simulated by the interpretation of non-logical symbols at worlds $w^{n}$.
More precisely, we define, by recursion, an infinite sequence $W^{0},W^{1},\ldots$ of subsets of $W$ and the interpretation of non-logical symbols at worlds from those sets in such a way that the interpretation of all non-logical symbols at all worlds from $W^{i}$, for a fixed $i$, is the same.
Define $W^{0}=\{w^{0}\}$ and,
$W^{i+1}=\bigl\{w\in W\mid\exists u\in W^{i}\ \exists v\in W\,(uR_{1}vR_{2}w~\&~u\neq v~\&~v\neq w)\bigr\},\qquad\text{for all }i\in\mathbb{N}.$
Then, define the interpretation at worlds $w$ from each $W^{i}$ inductively as follows:
  • $I_{w}(p)=1$;
  • $I_{w}(q_{j})=1$ iff $q^{i}=q_{j}$;
  • $I_{w^{0}}(V_{1})=I_{w^{0}}(V_{2})=\emptyset$;
  • $I_{w}(V_{1})$ and $I_{w}(V_{2})$, for $w\in W^{i}$ with $i>0$, depend on the instruction $I^{i-1}$ executed at step (i1)(i-1) of $C_{0}$ (below, $v$ is an arbitrary world from $W^{i-1}$; the choice does not matter since the interpretation of symbols is uniform in $W^{i-1}$):
    • if $I^{i-1}=+(r_{k},q_{j})$, then $I_{w}(V_{k})=I_{v}(V_{k})\cup\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{i-1}=-(r_{k},q_{j},q_{l})$ and $I_{v}(V_{k})\neq\emptyset$, then $I_{w}(V_{k})=I_{v}(V_{k})\setminus\{I_{v}(c)\}$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
    • if $I^{i-1}=-(r_{k},q_{l},q_{j})$ and $I_{v}(V_{k})=\emptyset$, then $I_{w}(V_{k})=I_{v}(V_{k})$ and $I_{w}(V_{3-k})=I_{v}(V_{3-k})$;
  • $I_{w}(c)$ depends on the instruction $I^{i}$ executed at step $i$ of $C_{0}$:
    • if $I^{i}=+(r_{k},q_{j})$ or $I^{i}=-(r_{k},q_{j},q_{l})$ and $I_{w}(V_{k})=\emptyset$, then $I_{w}(c)$ is any fixed $I_{w}(c)\notin I_{w}(V_{k})$;
    • if $I^{i}=-(r_{k},q_{l},q_{j})$ and $I_{w}(V_{k})\neq\emptyset$, then $I_{w}(c)$ is any fixed $I_{w}(c)\in I_{w}(V_{k})$.
Lastly, for $w\in W\setminus\bigcup\nolimits_{i\in\mathbb{N}}W^{i}$, we put $I_{w}(p)=0$ and $I_{w}(q_{i})=0$, for every $0\leq i\leq N$; the interpretation of other non-logical symbols at such worlds is immaterial. Then, $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$ and $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$.
$(\Leftarrow)$ Suppose there is a model $\mathfrak{M}=((W,R_{1},R_{2}),\langle\Delta,I_{w}\rangle_{w\in W})$ such that $\mathfrak{M}\models\varphi_{Q}\wedge\varphi_{I}$ and $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$, for some $w^{0}\in W$. We define an infinite sequence $(q^{0},v^{0}_{1},v^{0}_{2}),\,(q^{1},v^{1}_{1},v^{1}_{2}),\,(q^{2},v^{2}_{1},v^{2}_{2}),\,\ldots$ of configurations of a Minsky machine and show that this sequence is the computation of $M$ on $(0,0)$. First, let
  • $q^{0}$ be a uniquely determined $q_{i}$, for $0\leq i\leq N$, such that $I_{w^{0}}(q_{i})=1$;
  • $v^{0}_{k}=|I_{w^{0}}(V_{k})|$, for $i=1,2$.
Since $\mathfrak{M},w^{0}\models\varphi_{\mathit{init}}$ and $\mathfrak{M}\models\varphi_{Q}$, it follows that $(q^{0},v^{0}_{1},v^{0}_{2})$ is the initial configuration on $(0,0)$. From there, we can follow formulas $\varphi_{i}$ to construct the non-terminating computation of $M$ on $(0,0)$. ∎
We conclude this section with two remarks. First, note that neither $\mathcal{D}$ nor $\mathcal{D}_{\textup{fin}}$ are propositionally modally definable classes of frames. In fact, as $\mathbf{Diff}$ is axiomatised by $p\rightarrow\Box\Diamond p$ and $\Diamond p\rightarrow\Box(p\vee\Diamond p)$, any frame $(W,E)$ with an equivalence relation $E$ validates Diff (in other words, $\mathbf{Diff}\subseteq\mathbf{S5}$). The question arises whether our non‑preservation result continues to hold when one restricts the components to propositionally modally definable frame classes. We conjecture that it does, but at present we only have a proof for expanding domain semantics and under the additional assumption that polymodal logics are permitted as components. To demonstrate this, let $\mathcal{C}_{\text{diff}}$ be the class of all frames validating Diff and $\mathcal{C}_{\text{ftime}}$ the class of all frames validating the bimodal logic $\textbf{LTL}_{f}$ of all frames $(W,S,R)$ with $R$ a finite strict linear order and $S$ its successor relation. $\textbf{LTL}_{f}$ is axiomatised by taking the fusion of GL.3 and $\textbf{Alt}_{1}$ and adding the standard induction axioms [16].
Theorem 4.5.
$\mathop{\textup{Log}_{\textit{xd}}^{=}}\mathcal{C}_{\textup{diff}}$ and $\mathop{\textup{Log}_{\textit{xd}}^{=}}\mathcal{C}_{\textup{ftime}}$ are decidable but $\mathop{\textup{Log}_{\textit{xd}}^{=}}(\mathcal{C}_{\textup{diff}}\otimes\mathcal{C}_{\textup{diff}}\otimes\mathcal{C}_{\textup{ftime}})$ is undecidable.
Proof.
Decidability of $\mathop{\textup{Log}_{\textit{xd}}^{=}}\mathcal{C}_{\text{diff}}$ is shown in [20]. For decidability (and Ackermann-hardness) of $\mathop{\textup{Log}_{\textit{xd}}^{=}}\mathcal{C}_{\text{ftime}}$, see [18, 13]. Undecidability of $\mathop{\textup{Log}_{\textit{xd}}^{=}}(\mathcal{C}_{\text{diff}}\otimes\mathcal{C}_{\text{diff}}\otimes\mathcal{C}_{\text{ftime}})$ can be shown by adapting the reduction of Diophantine equations: first, observe that one can enforce that a frame in $\mathcal{C}_{\text{diff}}$ is actually in $\mathcal{D}$ using the formula
$\Box^{+}\forall x\,\bigl((x=c)\rightarrow\Box(x\neq c)\bigr),$
which can only be satisfied if all worlds in $W$ are irreflexive. The fusion with $\mathcal{C}_{\text{ftime}}$ is used to enforce a finite first-order domain $\Delta$ for worlds in $W$. This can be achieved as follows: for another fresh constant $c^{\prime}$ and $\Box_{L}$ the modal operator interpreted by the strict order in $\mathcal{C}_{\text{ftime}}$, if the sentence $\forall x\,\Diamond_{L}^{+}(x=c^{\prime})$ is true at $w_{0}$, then the domain $\Delta_{w_{0}}$ at $w_{0}$ cannot be larger than the length of the finite strict order seen from $w_{0}$. It is now straightforward to adapt the remaining steps of the undecidability proof. ∎
We conjecture that using polymodal logics here is not essential. For instance, the standard reduction of polymodal logics to logics with one modality might be applicable [22].
Second, the undecidability proof can be adapted to resolve an interesting open problem concerning monodic fragments of first‑order modal logics [23, 1]. A key ingredient of the encoding of Diophantine equations is that we can enforce equal cardinality $|P|_{w}=|Q|_{w}$ of the extensions of predicates $P$ and $Q$ at world $w$. This can also be achieved directly using a bijective function in the extension of one-variable modal logic to, for instance, the monodic fragment of modal first-order logic over the two-variable fragment of $\mathbf{FO}$ with counting. Hence, it can be shown that this monodic fragment is undecidable over the class $\mathcal{D}_{\text{fin}}$. This shows that decidability does not always transfer from one-variable modal logic with equality to monodic fragments based on decidable fragments of first-order logic (recall that the two-variable fragment with counting is decidable [27]). We remind that reader that sufficient conditions for this kind of preservation from one-variable modal logics to monodic fragments are shown in [14].
Theorem 4.6.
$\mathop{\textup{Log}_{\textit{cd}}^{=}}\mathcal{D}_{\textup{fin}}$ is decidable, but the monodic fragment of the two-variable fragment with counting over $\mathcal{D}_{\textup{fin}}$ is undecidable.
5 Fusions of Propositional Modal Logics Sharing an S5 Modality
In this section, we generalise fusions of one‑variable first‑order modal logics without equality to fusions of propositional modal logics sharing an S5 modality. We first establish a transfer result for logics admitting models homogeneous for the shared modality and then apply it to (semi)commutators.
Consider a propositional modal language with S5 operator $\Box_{E}$ interpreted in Kripke models by the equivalence relation $E$. The $E$-equivalence class $w]_{E}$ of $w\in W$ is defined as usual by setting $w]_{E}=E(w)$. Let $\kappa$ be an infinite cardinal. Then $\mathfrak{M}$ is called $(E,\kappa)$-homogeneous if, for every $E$-equivalence class $U$ in $\mathfrak{F}$ and every modal formula $\varphi$, $|\{w\in U\mid\mathfrak{M},w\models\varphi\}|\in\{0,\kappa\}$.
Definition 5.1.
A propositional modal logic $L$ with S5 operator $\Box_{E}$ admits $E$-homogeneous models (for global consequence) if there exists a cardinal $\kappa$ such that, for all $\kappa^{\prime}\geq\kappa$, whenever $\varphi\notin L$ (or, respectively, $\varphi\not\vdash^{\ast}_{L}\psi$), there exists an $(E,\kappa^{\prime})$-homogeneous model $\mathfrak{M}$ based on an $L$-frame with $\mathfrak{M}\not\models\varphi$ (or, respectively, with $\mathfrak{M}\models\varphi$ and $\mathfrak{M}\not\models\psi$).
Note that modal logics that admit homogeneous models (for global consequence) are (globally) Kripke complete, by definition. To formulate our main results, assume for simplicity that $L_{i}$ is a bimodal logic in the propositional modal language $\mathcal{L}_{i,E}$ with modal operators $\Box_{i}$ and $\Box_{E}$, for $i=1,2$. Then $L_{1}\otimes_{E}L_{2}$ denotes the smallest modal logic containing $L_{1}\cup L_{2}$ in the language $\mathcal{L}_{1,2,E}$ with modal operators $\Box_{1},\Box_{2}$, and $\Box_{E}$.
Theorem 5.2.
Let $L_{1}$ and $L_{2}$ admit $E$-homogeneous models (for global consequence). Then
  • $L_{1}\otimes_{E}L_{2}$ is (globally) Kripke complete;
  • $L_{1}\otimes_{E}L_{2}$ is (globally) decidable if $L_{1}$ and $L_{2}$ are both (globally) decidable.
Modulo replacing $\forall x$ with $\Box_{E}$, we use the same notation as in Section 3. For instance, if $\Phi$ is a finite set of $\mathcal{L}_{1,2,E}$-formulas closed under subformulas, then a $\Phi$-type is a maximal Boolean-consistent subset of $\{\psi,\neg\psi\mid\psi\in\Phi\}$, and a $\Phi$-quasistate is a non-empty set of $\Phi$-types satisfying the $\Box_{E}$-counterpart of (qs). We set
$\hat{\boldsymbol{q}}\ \ =\ \ \Box_{E}\bigvee_{t\in\boldsymbol{q}}t\ \ \land\ \ \bigwedge_{t\in\boldsymbol{q}}\Diamond_{E}t,$
where, for simplicity, we use $t$ to denote the conjunction of all formulas in $\Phi$-type $t$. Again, $\hat{\boldsymbol{Q}}$ denotes $\bigvee\{\hat{\boldsymbol{q}}\mid\boldsymbol{q}\in\boldsymbol{Q}\}$ for a set $\boldsymbol{Q}$ of $\Phi$-quasistates. We also define surrogates in the obvious way: for each formula of the form $\Box_{i}\psi$, we introduce a fresh proposition letter $p_{\Box_{i}\psi}$ and call it the surrogate of $\Box_{i}\psi$. As before, for any $\mathcal{L}_{1,2,E}$-formula $\chi$, we denote by $\chi^{\mathbf{1}}$ the $\mathcal{L}_{1,E}$-formula obtained from $\chi$ by replacing $\Box_{2}\psi$ not in the scope of another $\Box_{2}$ with its surrogate; $\chi^{\mathbf{2}}$ is defined symmetrically.
To establish the claimed results in the global case, it suffices to show the following lemma.
Lemma 5.3.
Let $\varphi,\psi$ be $\mathcal{L}_{1,2,E}$-formulas. Denote $\Phi=\mathop{\textit{sub}}\varphi\cup\mathop{\textit{sub}}\psi$. Then the following are equivalent:
  • (i)
    $\varphi\not\vdash_{L_{1}\otimes_{E}L_{2}}^{*}\psi$;
  • (ii)
    there is a model $\mathfrak{M}$ based on an $L_{1}\otimes_{E}L_{2}$-frame such that $\mathfrak{M}\models\varphi$ but $\mathfrak{M}\not\models\psi$;
  • (iii)
    there is a set $\boldsymbol{Q}$ of $\Phi$-quasistates such that
    • (C1)
      $\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}\not\vdash_{L_{1}}^{*}\psi^{\mathbf{1}}$,
    • (C2)
      $\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}\not\vdash_{L_{1}}^{*}\neg\hat{\boldsymbol{q}}_{i}^{\mathbf{1}}$, for each $\boldsymbol{q}_{i}\in\boldsymbol{Q}$,
    • (C3)
      $\smash{\hat{\boldsymbol{Q}}}^{\mathbf{2}}\not\vdash_{L_{2}}^{*}\neg\hat{\boldsymbol{q}}_{i}^{\mathbf{2}}$, for each $\boldsymbol{q}_{i}\in\boldsymbol{Q}$.
Proof.
Implications $(\textit{ii})\Rightarrow(\textit{i})$ and $(\textit{i})\Rightarrow(\textit{iii})$ are similar to the one-variable case and omitted.
$(\textit{iii})\Rightarrow(\textit{ii})$ Let $\boldsymbol{Q}$ be a set of $\Phi$-quasistates such that (C1)(C3) hold.
Choose an infinite $\kappa$ such that if $\chi\not\models_{L_{i}}^{\ast}\chi^{\prime}$, then for every $\kappa^{\prime}\geq\kappa$ there exists an $(E,\kappa^{\prime})$-homogeneous model $\mathfrak{M}_{i}$ based on an $L_{i}$-frame with $\mathfrak{M}_{i}\models\chi$ and $\mathfrak{M}_{i}\not\models\chi^{\prime}$, for $i=1,2$. By choosing $\kappa^{\prime}\geq\kappa$ sufficiently large and taking appropriate disjoint unions of models, it follows from (C1)(C3) that, for each $i=1,2$, we can find an $(E_{i},\kappa^{\prime})$-homogeneous model $\mathfrak{M}_{i}=(\mathfrak{F}_{i},V_{i})$ with $\mathfrak{F}_{i}=(W_{i},R_{i},E_{i})$ such that $\mathfrak{F}_{i}\models L_{i}$ and the following conditions hold:
  1. 1.
    $\mathfrak{M}_{1}\models\varphi^{\mathbf{1}}\land\smash{\hat{\boldsymbol{Q}}}^{\mathbf{1}}$, $\mathfrak{M}_{1}\not\models\psi^{\mathbf{1}}$, and $|\{[w]_{E_{1}}\mid\mathfrak{M}_{1},w\models\hat{\boldsymbol{q}}^{\mathbf{1}}\}|=\kappa^{\prime}$ for all $\boldsymbol{q}\in\boldsymbol{Q}$;
  2. 2.
    $\mathfrak{M}_{2}\models\smash{\hat{\boldsymbol{Q}}}^{\mathbf{2}}$ and $|\{[w]_{E_{2}}\mid\mathfrak{M}_{2},w\models\hat{\boldsymbol{q}}^{\mathbf{2}}\}|=\kappa^{\prime}$ for all $\boldsymbol{q}\in\boldsymbol{Q}$.
It follows that we find a bijection $F\colon W_{1}\rightarrow W_{2}$ such that, for all $w\in W_{1}$:
  1. 1.
    there is a (uniquely determined) $\Phi$-type $t$ such that $\mathfrak{M}_{1},w\models t^{\mathbf{1}}$ and $\mathfrak{M}_{2},f(w)\models t^{\mathbf{2}}$;
  2. 2.
    $f(w)]_{E_{2}}=\{f(w^{\prime})\mid w^{\prime}\in[w]_{E_{1}}\}$.
Now construct a new model $\mathfrak{M}=(\mathfrak{F},V)$ by taking $\mathfrak{M}_{1}$, interpreting $\Box_{2}$ by $f^{-1}[R_{2}]=\{(w,w^{\prime})\in W_{1}\mid(f(w),f(w^{\prime}))\in R_{2}\}$, and setting $V(p)=f^{-1}(V_{2}(p))$ for all 2-surrogate variables $p$. Then, for $\mathfrak{F}=(W_{1},R_{1},f^{-1}[R_{2}],E_{1})$, we have $\mathfrak{F}\models L_{1}\otimes_{E}L_{2}$ by definition, and one can show by induction for all formulas $\chi\in\Phi$ and $w\in W_{1}$:
$\mathfrak{M},w\models{\chi}\qquad\text{ iff }\qquad\mathfrak{M}_{1},w\models\chi^{\mathbf{1}}\qquad\text{ iff }\qquad\mathfrak{M}_{2},f(w)\models\chi^{\mathbf{2}}.$
It follows that $\mathfrak{M}\models\varphi$ and $\mathfrak{M}\not\models\psi$, as required. ∎
We next come to the local case. We use again the notation introduced for one-variable modal logics, with a few adaptations. For $\varphi$ an $\mathcal{L}_{1,2,E}$-formula, we let $\Theta^{\mathbf{1}}(\varphi)$ be the set of propositional variables in $\varphi$ and all $\chi\in\mathop{\textit{sub}}\Box_{2}\psi$ with $\Box_{2}\psi\in\mathop{\textit{sub}}\varphi$. An important difference is the relevant distance between nodes in frames. We define $\textit{dist}_{R,E}(w,w^{\prime})$ as the length of the shortest $S$-path from $w$ to $w^{\prime}$ with $S=E\circ R\circ E$. For $E$-equivalence classes $T,T^{\prime}$ we set $\textit{dist}_{R,E}(T,T^{\prime})=\textit{dist}_{R,E}(w,w^{\prime})$ with $w\in T$ and $w^{\prime}\in T^{\prime}$. We also set, inductively, $\Box_{i,E}^{1}\chi=\Box_{E}\Box_{i}\Box_{E}\chi$, $\Box_{i,E}^{m+1}\chi=\Box_{E}\Box_{i}\Box_{E}\Box_{i,E}^{m}\chi$ and similar for $\Box_{i,E}^{\leq m}\chi$, for $i=1,2$.
Lemma 5.4.
For all $\mathcal{L}_{1,2,E}$-formulas $\varphi$ and $i=1,2$ with $\textit{adp}(\varphi)=\textit{adp}^{\mathbf{i}}{(\varphi)}$, the following are equivalent:
  • $\not\vdash_{L_{1}\otimes_{E}L_{2}}\varphi$;
  • there is a model $\mathfrak{M}$ based on an $L_{1}\otimes_{E}L_{2}$-frame such that $\mathfrak{M}\not\models\varphi$;
  • $\not\vdash_{L_{i}}\displaystyle\Box_{i,E}^{\leq\textit{md}^{\mathbf{i}}(\varphi)}(\hat{\boldsymbol{Q}}_{i}(\varphi))^{\mathbf{i}}\to\varphi^{\mathbf{i}}$, where $\boldsymbol{Q}_{i}(\varphi)$ is the set of all $\Theta^{\mathbf{i}}(\varphi)$-quasistates $\boldsymbol{q}$ such that $\not\vdash_{L_{1}\otimes_{E}L_{2}}\neg\hat{\boldsymbol{q}}$.
Proof.
Let $\kappa$ be an infinite cardinal such that, if $\chi\not\in L_{i}$, then for every $\kappa^{\prime}\geq\kappa$ there exists an $(E,\kappa^{\prime})$-homogeneous model $\mathfrak{M}_{i}$ based on an $L_{i}$-frame with $\mathfrak{M}_{i}\not\models\chi$, for $i=1,2$. Let $\Phi$ be a set of $\mathcal{L}_{1,2,E}$-formulas closed under subformulas. For $i=1,2$ and $h\leq\textit{md}^{\mathbf{i}}(\Phi)$, let $\Gamma_{h}^{\mathbf{i}}(\Phi)=\{\psi\in\Phi\mid\textit{md}^{\mathbf{i}}(\psi)\leq h\}$. A tapered $(i,\kappa)$-quasimodel $\mathfrak{Q}$ for $\Phi$ is a tuple of the form $((W,R_{i},E),T_{0},\tau)$, where
  • $(W,R_{i},E)\models L_{i}$,
  • $T_{0}\subseteq W$ is an $E$-equivalence class, and
  • $\tau$ is a tapered mapping associating with each $w\in W$ a $\Gamma^{\mathbf{i}}_{h(w)}(\Phi)$-type $\tau(w)$, where $h(w)=\textit{md}^{\mathbf{i}}(\Phi)-\textit{dist}_{R_{i},E}(T_{0},[w]_{E})$,
such that the following conditions hold:
  1. 1.
    for every $E$-equivalence class $T$ and finite $\Xi$ of $\mathcal{L}_{1,2,E}$-formulas, $|\{w\in T\mid\tau(w)=\Xi\}|\in\{0,\kappa\}$;
  2. 2.
    $\tau$ is $i$-coherent at every $w\in W$: $\Box_{i}\psi\in\tau(w)$ implies $\psi\in\tau(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$;
  3. 3.
    $\tau$ is $i$-saturated at every $w\in W$: $\neg\Box_{i}\psi\not\in\tau(w)$ whenever $\psi\in\tau(w^{\prime})$, for all $w^{\prime}\in R_{i}(w)$;
  4. 4.
    $\tau$ is $E$-coherent at every $w\in W$: $\Box_{E}\psi\in\tau(w)$ implies $\psi\in\tau(w^{\prime})$, for all $w^{\prime}\in E(w)$;
  5. 5.
    $\tau$ is $E$-saturated at every $w\in W$: $\neg\Box_{E}\psi\not\in\tau(w)$ whenever $\psi\in\tau(w^{\prime})$, for all $w^{\prime}\in E(w)$.
We now inductively define a tapered $(i,\kappa)$-cactus $\mathfrak{C}$ for $\Phi$ as a tuple of the form $((W,R_{1},R_{2},E),T_{0},\tau)$, where $(W,R_{i},E)\models L_{i}$, $T_{0}\subseteq W$ is an $E$-equivalence class, no node in $T_{0}$ is in the domain or range of $R_{3-i}$, $(W\setminus T_{0},R_{3-i},E\setminus T_{0}\times T_{0})\models L_{3-i}$, and $\tau$ is a tapered mapping such that, for every set $\Xi$ of formulas and $E$-equivalence class $T$, $|\{w\in T\mid\tau(w)=\Xi\}|\in\{0,\kappa\}$ and which is $1$-, 22- and $E$-coherent, and $i$- and $E$-saturated on $W$ but (3i)(3-i)-saturated only on $W\setminus T_{0}$.
For the base case, one can easily show that a ‘degenerate’ tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),T_{0},\tau)$ for $\emptyset$ always exists. Note that, as $\Phi=\emptyset$, we have $\tau(w)=\emptyset$ for all $w\in W$ and $|T_{0}|=\kappa$.
For the inductive case, $\Phi\neq\emptyset$. Suppose $\mathfrak{Q}=((W^{0},R_{1}^{0},E^{0}),T^{0},\tau^{0})$ is a tapered $(1,\kappa)$-quasimodel for $\Phi$. Let $\mathcal{T}$ denote the set of $E^{0}$-equivalence classes $T\subseteq W^{0}$ with $T\not=T^{0}$. For each $T\in\mathcal{T}$, denote
$\Phi^{T}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(T)}(\Phi),\quad\text{ where }h(T)=\textit{md}^{\mathbf{1}}(\Phi)-\textit{dist}_{R_{1}^{0},E}(T_{0},T).$
For each $T\in\mathcal{T}$, take a tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}=((W^{T},R_{1}^{T},R_{2}^{T},E^{T}),T,\tau^{T})$ for $\Phi^{T}$ that is compatible with $\mathfrak{Q}$ in the sense that
$W^{0}\cap W^{T}=T\quad\text{ and }\quad\tau^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}=\tau^{T}(v)\text{ for all $v\in T$}.$
We also assume that the $W^{T}$ are disjoint. Then $\mathfrak{C}=((W,R_{1},R_{2},E),T^{0},\tau)$ is a tapered $(1,\kappa)$-cactus for $\Phi$ if it is obtained by grafting, for each $T\in\mathcal{T}$, the tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ onto $T$ in $\mathfrak{Q}$ as follows:
$W\ \ =\ \ T^{0}\ \cup\ \bigcup\nolimits_{T\in\mathcal{T}}W^{T},\qquad R_{1}\ \ =\ \ R_{1}^{0}\ \cup\ \bigcup\nolimits_{T\in\mathcal{T}}R_{1}^{T},\qquad R_{2}\ \ =\ \ \bigcup\nolimits_{T\in\mathcal{T}}R_{2}^{T}\\ E\ \ =\ \ E^{0}\ \cup\ \bigcup\nolimits_{T\in\mathcal{T}}E^{T},\qquad\tau\ \ =\ \ \tau^{0}\ \cup\ \bigcup\nolimits_{T\in\mathcal{T}}\tau^{T}.$
A fully grafted tapered $(i,\kappa)$-cactus $((W,R_{1},R_{2},E),\tau)$ for $\Phi$ results from grafting a compatible tapered $((3-i),\kappa)$-cactus $\mathfrak{C}^{T^{0}}$ for $\Phi^{T^{0}}=\Theta^{\mathbf{i}}(\Phi)$ onto $T^{0}$ in a tapered $(i,\kappa)$-cactus $((W^{0},R_{1}^{0},R_{2}^{0},E^{0}),T^{0},\tau^{0})$ for $\Phi$. Observe that the resulting mapping $\tau$ is $1$-, 22- and $E$-coherent and $1$-, 22- and $E$-saturated on $W$.
We show by induction on $\textit{adp}(\varphi)$, for both $i=1,2$, that, if $\textit{adp}(\varphi)=\textit{adp}^{\mathbf{i}}{(\varphi)}$, then the following are equivalent:
  • (i)
    $\not\vdash_{L_{1}\otimes_{E}L_{2}}\varphi$;
  • (ii)
    there is a model $\mathfrak{M}$ based on an $L_{1}\otimes_{E}L_{2}$-frame such that $\mathfrak{M}\not\models\varphi$;
  • (iii)
    there is a (fully grafted) tapered $(i,\kappa)$-cactus for $\mathop{\textit{sub}}\varphi$ with $\varphi\notin\tau(w_{0})$ for some $T_{0}$ and $w_{0}\in T_{0}$;
  • (iv)
    $\not\vdash_{L_{i}}\displaystyle\Box_{i,E}^{\leq\textit{md}^{\mathbf{i}}(\varphi)}(\hat{\boldsymbol{Q}}_{i}(\varphi))^{\mathbf{i}}\to\varphi^{\mathbf{i}}$.
The basis of induction, $\textit{adp}(\varphi)=0$, is trivial as the formula belongs to $\mathcal{L}_{1,E}$ or $\mathcal{L}_{2,E}$ (or both, if it has no modalities). For the inductive step, consider an $\mathcal{L}_{1,2,E}$-formula $\varphi$ and suppose that $\textit{adp}(\varphi)=k>0$. Assume that $\textit{adp}^{\mathbf{1}}(\varphi)=k$ (the other case is symmetrical) and that the claim has been shown for all formulas $\psi$ with $\textit{adp}^{\mathbf{2}}(\psi)<k$, which, by ($\dagger$3.2), includes all $\psi\in\Theta^{\mathbf{1}}(\varphi)$.
Implication $(\textit{ii})\Rightarrow(\textit{i})$ follows from soundness of $\vdash_{L_{1}\otimes_{E}L_{2}}$.
$(iv)\Rightarrow(iii)$ By definition of $\kappa$, there is an $(E,\kappa)$-homogeneous model $\mathfrak{M}^{0}$ based on an $L_{1}$-frame $(W^{0},R^{0}_{1},E^{0})$ such that $\mathfrak{M}^{0},w_{0}\models\Box_{1,E}^{\leq\textit{md}^{\mathbf{1}}(\varphi)}(\hat{\boldsymbol{Q}}_{1}(\varphi))^{\mathbf{1}}$ and $\mathfrak{M}^{0},w_{0}\not\models\varphi^{\mathbf{1}}$ for some $w_{0}\in W^{0}$. Let $T^{0}=[w_{0}]_{E^{0}}$ and $\Phi=\mathop{\textit{sub}}\varphi$. For each $E^{0}$-equivalence class $T\subseteq W^{0}$, let $h(T)=\textit{md}^{\mathbf{1}}(\Phi)-\textit{dist}_{R^{0}_{1},E^{0}}(T^{0},T)$. Then, for each $v\in W^{0}$ in each $E^{0}$-equivalence class $T=[v]_{E^{0}}$, we take the $\Gamma^{\mathbf{1}}_{h(T)}(\Phi)$-type $t_{v}$ of $v$ such that $\mathfrak{M},v\models t_{v}^{\mathbf{1}}$. Define the tapered mapping $\tau^{0}$ through $(W^{0},R^{0}_{1},E^{0})$ by setting $\tau^{0}(v)=t_{v}$ for $v\in W^{0}$. It can be seen $\mathfrak{Q}=((W^{0},R^{0}_{1},E^{0}),T^{0},\tau^{0})$ is a tapered $(1,\kappa)$-quasimodel for $\Phi$ such that $\varphi\notin t_{w_{0}}$.
For each $E^{0}$-equivalence class $T$ with $h(T)\geq 0$, consider the $\Theta^{1}(\Phi)$-quasistate $\boldsymbol{q}^{T}\in\boldsymbol{Q}_{1}(\varphi)$ such that $\mathfrak{M}^{0},v\models(\hat{\boldsymbol{q}}^{T})^{\mathbf{1}}$ for all (equivalently, some) $v\in T$. Observe that $\boldsymbol{q}^{T}$ may contain formulas $\chi$ with $\textit{md}^{\mathbf{1}}{(\chi)}>h(T)$. Let $\Phi^{T}=\Theta^{\mathbf{1}}(\Phi)\cap\Gamma^{\mathbf{1}}_{h(T)}(\Phi)$ and consider the restriction $\boldsymbol{q}_{0}^{T}=\{t|_{\smash{\Gamma^{\mathbf{1}}_{h(T)}(\Phi)}}\mid t\in\boldsymbol{q}^{T}\}$, which is a $\Phi^{T}$-quasistate. Denote $\smash{\hat{\boldsymbol{q}}^{T}_{0}}$ by $\varphi^{T}$. Since $\textit{adp}^{\mathbf{2}}(\varphi^{T})<k$, we apply our induction hypothesis to the $\mathcal{L}_{1,2,E}$-formula $\varphi^{T}$ and obtain a tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}=((W^{T},R_{1}^{T},R_{2}^{T},E^{T}),T,\tau^{T})$ for $\mathop{\textit{sub}}\varphi^{T}$ (equivalently, for $\Phi^{T}$) such that $\tau^{0}(v)|_{\Theta^{\mathbf{1}}(\Phi)}=\tau^{T}(v)$ for all $v\in T$. At each $E^{0}$-equivalence class $T$ with $h(T)<0$, we take a degenerate tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ for $\emptyset$. Next, for each $T\neq T_{0}$, we graft the tapered $(2,\kappa)$-cactus $\mathfrak{C}^{T}$ onto $T$ in the $(1,\kappa)$-quasimodel $\mathfrak{Q}$ to obtain a tapered $(1,\kappa)$-cactus $\mathfrak{C}$. A fully grafted tapered $(1,\kappa)$-cactus is obtained by grafting, in addition, $\mathfrak{C}^{T^{0}}$ onto $T^{0}$ in $\mathfrak{C}$.
$(iii)\Rightarrow(ii)$ Let $\mathfrak{C}=((W,R_{1},R_{2},E),\tau)$ be a fully grafted tapered $(i,\kappa)$-cactus for $\mathop{\textit{sub}}\varphi$ with $\varphi\notin\tau(w_{0})$, for some $w_{0}\in T_{0}$. We construct a Kripke model $\mathfrak{M}$ based on $(W,R_{1},R_{2},E)$ by setting $\mathfrak{M},w\models p$ iff $p\in\tau(w)$, for all proposition letters $p$ and $w\in W$. By induction on the structure of a formula $\chi$, one can show for all $w\in W$, $\chi\in\{\psi,\neg\psi\mid\psi\in\tau(w)\}$, and both $i=1,2$,
$\mathfrak{M},w\models\chi\qquad\text{ iff }\qquad\chi^{\mathbf{i}}\in(\tau(w))^{\mathbf{i}}.$
The basis of induction is by definition. The Boolean connectives are trivial, the inductive step for modal operator $\Box_{E}$ follows from the condition that $\tau$ is $E$-coherent and $E$-saturated, and the inductive step for $\Box_{i}$, $i=1,2$, follows from the condition that $\tau$ is $i$-coherent and $i$-saturated. Thus, $\mathfrak{M},w_{0}\not\models\varphi$.
$(i)\Rightarrow(iv)$ is trivial. ∎
We next show that the results above apply to Kripke complete (semi)commutators with S5.
Lemma 5.5.
If a logic of the form $L,\mathbf{S5}]^{\mathit{xd}}$ or $L,\mathbf{S5}]^{\textit{cd}}$ is (globally) Kripke complete, then it admits $E$-homogeneous models (for global consequence).
Proof.
We give the proof for the local case, the proof in the global case is the same. Assume first that $L,\textbf{S5}]^{\textit{xd}}$ is Kripke complete. Let $\kappa$ be any infinite cardinal such that, for every $\varphi\not\in[L,\textbf{S5}]^{\textit{xd}}$, there exists a model refuting $\varphi$ based on an $L,\textbf{S5}]^{\textit{xd}}-$frame of size at most $\kappa$. We show that $\kappa$ is as required. Assume $\kappa^{\prime}\geq\kappa$ and $\varphi\not\in[L,\textbf{S5}]^{\textit{xd}}$. By definition of $\kappa$, there exists a model $\mathfrak{M}=(\mathfrak{F},V)$ with $\mathfrak{F}=(W,R,E)$, $\mathfrak{M}\not\models\varphi$, $\mathfrak{F}\models[L,\textbf{S5}]^{\textit{xd}}$, and $W$ of size $\kappa$. Then define $\mathfrak{M}_{\kappa^{\prime}}=(\mathfrak{F}_{\kappa^{\prime}},V_{\kappa^{\prime}})$ with $\mathfrak{F}_{\kappa^{\prime}}=(W^{\prime},R^{\prime},E^{\prime})$ by setting
  • $W^{\prime}=W\times\kappa^{\prime}$;
  • $(w,n)R^{\prime}(w^{\prime},n^{\prime})$ if $wRw^{\prime}$ and $n=n^{\prime}$, for all $n,n^{\prime}\in\kappa^{\prime}$;
  • $(w,n)E^{\prime}(w^{\prime},n^{\prime})$ if $wEw^{\prime}$, for all $n,n^{\prime}\in\kappa^{\prime}$;
  • $V_{\kappa^{\prime}}(p)=V(p)\times\kappa^{\prime}$, for all proposition letters $p$.
Claim. $\mathfrak{F}_{\kappa^{\prime}}\models[L,\textbf{S5}]^{\textit{xd}}$, $\mathfrak{M}_{\kappa^{\prime}}\not\models\varphi$, and $\mathfrak{M}_{\kappa^{\prime}}$ is an $(E,\kappa^{\prime})$-homogeneous model.
Proof of Claim. By closure under disjoint unions, $(W^{\prime},R^{\prime})$ validates $L$. Also, $E^{\prime}$ is easily seen to be an equivalence relation. We next show that $(W^{\prime},R^{\prime},E^{\prime})$ satisfies $(\textit{lcom}_{R,E})$. Assume $(w,n)E^{\prime}(w^{\prime},m)$ and $(w^{\prime},m)R^{\prime}(w^{\prime\prime},k)$. By definition, $wEw^{\prime}$, $m=k$, and $w^{\prime}Rw^{\prime\prime}$. Then, as $(W,R,E)$ satisfies $(\textit{lcom}_{R,E})$, we have a $v^{\prime}$ with $wRv^{\prime}$ and $v^{\prime}Ew^{\prime\prime}$. But then $(w,n)R^{\prime}(v^{\prime},n)$ and $(v^{\prime},n)E^{\prime}(w^{\prime\prime},k)$, as required.
Let $\chi$ be any formula. Then one can easily show by induction that $\mathfrak{M},w\models\chi$ iff $\mathfrak{M}_{\kappa^{\prime}},(w,n)\models\chi$ for all $n\in\kappa^{\prime}$. It follows that $\mathfrak{M}_{\kappa^{\prime}}\not\models\varphi$ and $\mathfrak{M}_{\kappa^{\prime}}$ is an $(E,\kappa^{\prime})$-homogeneous model. This finishes the proof of the claim and also the proof for logics of the form $L,\textbf{S5}]^{\textit{xd}}$.
Now assume a logic of the form $L,\mathbf{S5}]^{\textit{cd}}$ is given and Kripke complete. The construction of $\mathfrak{M}_{\kappa^{\prime}}$ is the same, using a model $\mathfrak{M}=(\mathfrak{F},V)$ with $\mathfrak{F}$ also satisfying $(\textit{rcom}_{R,E})$. So it suffices to show that $\mathfrak{M}_{\kappa^{\prime}}$ then also satisfies $(\textit{rcom}_{R,E})$: assume $(w,n)R^{\prime}(w^{\prime},m)$ and $(w^{\prime},m)E^{\prime}(w^{\prime\prime},k)$. By definition, $n=m$, $wRw^{\prime}$, and $w^{\prime}Ew^{\prime\prime}$. Then, as $(W,R,E)$ satisfies $(\mathit{rcom}_{R,E})$, we have a $v^{\prime}$ with $wEv^{\prime}$ and $v^{\prime}Rw^{\prime\prime}$. But then $(w,n)E^{\prime}(v^{\prime},k)$ and $(v^{\prime},k)R^{\prime}(w^{\prime\prime},k)$, as required. ∎
We obtain the following transfer result for (semi)commutators.
Theorem 5.6.
Let $d\in\{\textit{xd},\textit{cd}\}$ and $L_{1}$ and $L_{2}$ be consistent propositional modal logics such that $L_{i},\textbf{S5}]^{d}$ is (globally) Kripke complete, for both $i=1,2$. Then
  • $L_{1},\textbf{S5}]^{d}\otimes_{E}[L_{2},\textbf{S5}]^{d}$ is (globally) Kripke complete;
  • $L_{1},\textbf{S5}]^{d}\otimes_{E}[L_{2},\textbf{S5}]^{d}$ is (globally) decidable iff $L_{1},\textbf{S5}]^{d}$ and $L_{2},\textbf{S5}]^{d}$ are (globally) decidable.
6 Discussion
We have established preservation and non‑preservation results for one‑variable first‑order modal logics and for related propositional modal logics. Many questions remain open:
  • The behaviour of fusions of full first‑order modal logic is still largely unexplored. A transfer theorem for canonicity is given in [30], but, for example, the transfer of completeness is open.
  • Monodic fragments extend one‑variable modal logics by allowing more expressive first‑order fragments while restricting modal operators to formulas with at most one free variable [23, 1]. We have shown that decidability does not transfer from the one‑variable fragment with equality to the monodic fragment over the two‑variable fragment of FO with counting. Positive preservation results are known [14], and it is straightforward to adapt our proofs to, for instance, the monodic fragments based on the two-variable fragment of $\mathbf{FO}$ without equality or description logics such as ALC, ALCI, and SHI with the universal role [3, 23]. However, a systematic study of fusions of monodic fragments is still missing.
  • It would be valuable to identify sufficient conditions under which Kripke completeness and decidability transfer to fusions of one‑variable modal logics with equality. We note that many standard polymodal fusions with equality such as polymodal $\mathbf{K}$ and $\mathbf{S5}$ are in fact decidable [1].
  • We have seen that, in many cases, completeness and decidability are preserved under fusions of propositional modal logics sharing an S5 operator. Can this be extended to other modal logics?
  • Our analysis focused on Kripke complete modal logics. Since first‑order modal logics are often incomplete under standard Kripke semantics, it would be interesting to obtain preservation results that do not presuppose Kripke completeness. Techniques from [34, 2] may be helpful here.
Appendix A Proof of Lemma 2.1
Let $L^{\prime}=[L,\textbf{S5}]^{\textit{xd}}$ and let $\textbf{FO}_{1}$ be the one-variable fragment of the classical first-order logic. It is easy to check that $\vdash_{L^{\prime}}\Diamond\Box_{E}p\to\Box_{E}\Diamond p(=\textit{chr}_{\Diamond})$, for each $\Diamond$ corresponding to $\Box\in M$. We also rely on the well-known [33] fact that $\textbf{FO}_{1}^{\ast}=\textbf{S5}$.
First, we prove that the translation $\cdot^{\ast}$ commutes with substitution: for all one-variable formulas $\varphi,\psi_{1},\ldots,\psi_{n}$,
$\begin{array}[]{lcl}(\varphi)^{\ast}\,[\psi_{1},\ldots,\psi_{n}/P_{1}(x),\ldots,P_{n}(x)]&=&\varphi^{\ast}\,[\psi_{1}^{\ast},\ldots,\psi_{n}^{\ast}/p_{1},\ldots,p_{n}].\end{array}$ ()
The proof is by induction on $\varphi$; we only consider the case $\varphi=\exists x\,\theta$. Let $\sigma=[\psi_{1},\ldots,\psi_{n}/P_{1}(x),\ldots,P_{n}(x)]$ and $\sigma^{\ast}=[\psi_{1}^{\ast},\ldots,\psi_{n}^{\ast}/p_{1},\ldots,p_{n}]$. Since, by IH, $(\theta\,\sigma)^{\ast}=\theta^{\ast}\,\sigma^{\ast}$, it follows that
$\begin{array}[]{ccccccccccccc}((\exists x\,\theta)\,\sigma)^{\ast}&=&(\exists x\,(\theta\,\sigma))^{\ast}&=&\Diamond_{E}(\theta\,\sigma)^{\ast}&=&\Diamond_{E}(\theta^{\ast}\,\sigma^{\ast})&=&(\Diamond_{E}\theta^{\ast})\,\sigma^{\ast}&=&(\exists x\,\theta)^{\ast}\,\sigma^{\ast}.\end{array}$
We now prove that $L^{\prime}\subseteq\mathcal{Q}L^{\ast}$. It should be clear that $L\subseteq\mathcal{Q}L^{\ast}$, and that $\textbf{S5}=\textbf{FO}_{1}^{\ast}\subseteq\mathcal{Q}L^{\ast}$. Since $\textit{CBF}_{\Box}$ belongs to every one-variable modal logic and $\textit{CBF}_{\Box}^{\ast}=(\textit{lcom}_{\Box})$, it follows that $(\textit{lcom}_{\Box})\in\mathcal{Q}L^{\ast}$. Now, the claim follows from ($\ast$) and the closure of $\mathcal{Q}L^{\ast}$ under MP, Necessitation, and Generalization.
We next prove that $\mathcal{Q}L^{\ast}\subseteq L^{\prime}$. Here, we use the fact that $\mathbf{FO}_{1}$ is obtained by adding to the classical propositional logic the axiom $\forall x\,P(x)\to P(x)$ and the rule $\psi\to\varphi(x)/\psi\to\forall x\,\varphi(x)$, where $x$ is not free in $\psi$. In view of ($\ast$), the only non-trivial part is to show that $\vdash_{L^{\prime}}(\psi\to\varphi)^{\ast}$ implies that $\vdash_{L^{\prime}}(\psi\to\forall x\,\varphi)^{\ast}$ provided $x$ is not free in $\psi$. To that end, we show that, if $\gamma$ is a propositional $(M\cup\{\Box_{E}\})$-formula and and $\beta$ is a propositional $(M\cup\{\Box_{E}\})$-formula where all occurrences of proposition letters occur within the scope of $\Box_{E}$, then $\vdash_{L^{\prime}}\beta\to\gamma$ implies that $\vdash_{L^{\prime}}\beta\to\Box_{E}\gamma$. Since $\Box_{E}$ is monotonic, it is enough to show that $\vdash_{L^{\prime}}\beta\to\Box_{E}\beta$, for all $\beta$ of the assumed form.
Due to our assumption about its form, we may rewrite $\beta$ with $\neg$, $\wedge$, $\vee$, $\Box_{E}$, $\Box$, and $\Diamond$ so that all negations occur either within the scope of $\Box_{E}$ or immediately precede $\Box_{E}$ (to that end, we push all the negations outside of any $\Box_{E}$ down until they reach some $\Box_{E}$). Thus, $\beta$ is built from formulas of the form $\Box_{E}\delta$ and $\neg\Box_{E}\delta$ using $\wedge$, $\vee$, $\Box$, and $\Diamond$. We proceed by induction on the structure of such $\beta$.
  • If $\beta=\Box_{E}\gamma$ or $\beta=\neg\Box_{E}\gamma$, then the claim holds because $\Box_{E}$ is an S5 modality, and therefore $\vdash_{L^{\prime}}\Box_{E}\gamma\to\Box_{E}\Box_{E}\gamma$ and $\vdash_{L^{\prime}}\neg\Box_{E}\gamma\to\Box_{E}\neg\Box_{E}\gamma$.
  • The cases of $\wedge$ and $\vee$ are straightforward.
  • If $\beta=\Box\gamma$ and $\vdash_{L^{\prime}}\gamma\to\Box_{E}\gamma$, then, by monotonicity of $\Box$ and $(\textit{lcom}_{\Box})$, $\vdash_{L^{\prime}}\Box\gamma\to\Box_{E}\Box\gamma$.
  • If $\beta=\Diamond\gamma$ and $\vdash_{L^{\prime}}\gamma\to\Box_{E}\gamma$, then, by monotonicity of $\Diamond$ and $(\textit{chr}_{\Box})$, $\vdash_{L^{\prime}}\Diamond\gamma\to\Box_{E}\Diamond\gamma$.
To prove the claim of the lemma for $L,\textbf{S5}]^{\textit{cd}}$, it suffices to recall that $\textit{BF}_{\Box}^{\ast}=(\textit{rcom}_{\Box})$.